### v2

# Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation

Abstract
Detecting cracks with pixel-level precision for key structures is a significant challenge, as existing methods struggle to effectively integrate local textures and pixel dependencies of cracks. Furthermore, these methods often possess numerous parameters and substantial computational requirements, complicating deployment on edge control devices. In this paper, we propose a staircase cascaded fusion crack segmentation network (CrackSCF) that generates high-quality crack segmentation maps using minimal computational resources. We constructed a staircase cascaded fusion module that effectively captures local patterns of cracks and long-range dependencies of pixels, and it can suppress background noise well. To reduce the computational resources required by the model, we introduced a lightweight convolution block, which replaces all convolution operations in the network, significantly reducing the required computation and parameters without affecting the network’s performance. To evaluate our method, we created a challenging benchmark dataset called TUT and conducted experiments on this dataset and five other public datasets. The experimental results indicate that our method offers significant advantages over existing methods, especially in handling background noise interference and detailed crack segmentation. The F1 and mIoU scores on the TUT dataset are 0.8382 and 0.8473, respectively, achieving state-of-the-art (SOTA) performance while requiring the least computational resources. The code and dataset is available at https://github.com/Karl1109/CrackSCF.
Index Terms:
Structural cracks segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
I Introduction
Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1, 2]. Early automated detection methods used traditional image processing techniques for crack extraction [3, 4, 5]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
Figure 1: Performance comparison of our proposed method with UCTNet [6], SFIAN [7], CT-crackseg [8], DTrCNet [9], Crackmer [10] and Simcrack [11] on the TUT dataset in terms of mIoUs, FLOPs and Params metrics.
In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as MST-Net [12] and FcaHRNet [13] have shown efficient feature extraction capabilities in structural crack segmentation tasks. The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations. In labeled images, crack labels are long, narrow structures of white pixels, while the background is large black areas. This challenges the model to distinguish between thin cracks and the background. Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
The success of VIT [14] and Swin Transformer [15] has demonstrated the advantages of Transformer [16], with their strong sequence processing capabilities, in handling visual tasks. Unlike CNNs, transformers can capture dependencies between any pixels when processing images [17, 18], making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [19], MFAFNet [20], and Crackmer [9], which integrate local texture information and long-range pixel dependencies through U-shaped connections or residual connections.
Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations, as follows:
  1. 1.
    The different information fusion strategies significantly affect the segmentation performance of CNN-Transformer networks. For instance, UCTNet [6] and CT-crackseg [8] use channel concatenation to fuse local patterns and long-range dependences information. However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise. Additionally, DTrCNet [9] and Crackmer [10] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance. As shown in Fig. 1(a), due to these shortcomings, these four methods fail to achieve the best performance.
  2. 2.
    Most networks, in pursuit of performance, overlook the significant parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices. As shown in Fig. 1(b) and Fig. 1(c), networks like SFIAN [7], Crackmer [10], and CT-crackseg[8] have very high computational FLOPs and parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
To address these issues, we propose a staircase cascaded fusion network that can generate high-quality segmentation maps with minimal computational resources while effectively suppressing background noise. Specifically, to better exploit the complementarity and correlation of feature maps without missing critical detail information, we introduce a staircase cascaded fusion module. This module utilizes the pixel attention mechanism, which finely captures local and long-range information, and the channel concatenation mechanism, which progressively integrates features from adjacent layers, to generate crack segmentation maps that comprehensively combine local texture information and pixel dependencies. Furthermore, to decrease the parameter count and computational burden, we introduce a lightweight convolution block called LRDS. This block initially increases and then reduces the channel count, and separates the spatial and channel aspects, thereby minimizing computational costs while maintaining performance.
To assess the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [21] and Crack500 [22] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively assesses the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
In conclusion, our main contributions are as follows:
  1. 1.
    We proposed a novel CrackSCF network for structural crack segmentation. By introducing a staircase cascaded fusion module and a lightweight convolution block, this network effectively distinguishes between background and crack pixels with minimal computational resources, resulting in high-quality segmentation maps.
  2. 2.
    We collected the TUT dataset, which includes structural cracks in eight different scenarios, providing a more comprehensive evaluation of the network’s performance.
  3. 3.
    We carried out comprehensive experiments on five public datasets and the TUT benchmark dataset, demonstrating that our proposed CrackSCF network outperforms existing SOTA methods.
The organization of the subsequent sections of this paper is as follows: In Section II, existing crack detection models and lightweight networks are reviewed. Section III details the components of the proposed network. Section IV delves into the experimental details and results. Finally, Section V provides a summary of the entire paper and highlights future research directions.
II Related Work
II-A Crack Segmentation Networks
Recently, a growing number of researchers have utilized CNNs to develop crack segmentation networks, achieving advanced performance. Al-Huda et al. [23] proposed the MDAUNet network, based on U-Net and dual attention modules, and designed a hybrid weighted segmentation loss function to address class imbalance issues. Cheng et al. [7] introduced the Selective Feature Fusion and Irregular-Aware Network (SFIAN) to effectively model irregular crack objects, resulting in clearer backgrounds on various crack images. Huang et al. [24] developed a crack detection network for masonry structures that combines thermal imaging and visible light image fusion with semantic segmentation, laying the foundation for the development of cross-modal methods for structural crack monitoring.
Although the aforementioned methods have demonstrated the advantages of CNNs in crack segmentation tasks, CNNs are essentially local operations and cannot effectively perceive a small number of crack pixels amidst a large background. Therefore, these methods often result in discontinuous crack pixels and are highly susceptible to background noise interference.
Figure 2: General Architecture Diagram of CrackSCF Network. Crack images are first input into the MFE, generating four feature maps. After enhancement, five feature maps with unified channel numbers are obtained. The LDE module processes these maps to acquire pixel sequences rich in long-range correlations. The SCFM module then processes these sequences and the four layers of feature maps from the MFE. Through four stages, resulting in the segmentation output.
Thanks to the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes. Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation. Chen et al. [25] proposed TransUNet, a novel network architecture that effectively integrates long-range information and local features through a U-Net-like structure. Qi et al. [26] developed an end-to-end model based on Vision Transformer (ViT) and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders. Feng et al. [27] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
The aforementioned methods, while demonstrating the advantages of CNN-Transformer networks in processing pixel sequences and suppressing background noise, have limitations due to using only one method or branch for feature fusion. This reduces their ability to effectively perceive long-range dependencies and local patterns, causing some complex regions’ local or long-range information to be overshadowed. Additionally, these methods excessively focus on performance, overlooking the substantial computational resources required for feature extraction and fusion, making them difficult to deploy on edge computing devices.
II-B Lightweight Networks
Deploying deep learning models on edge devices like smartphones and drones is trending. However, these devices have limited computing resources, making it challenging to use models with high parameter counts. Thus, compact model design for lightweight modifications in size and speed is crucial. For example, Howard et al. [28] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices. Ma et al. [29] introduced ShuffleNet, providing practical design principles for efficient network architecture. Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures. Zhang et al. [30] proposed a low-rank technique for neural networks that employs constrained optimization to to determine the optimal low-rank approximation. Swaminathan et al. [31] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
Recently, many researchers have combined model lightweighting techniques with crack segmentation models to enable real-time crack detection on edge devices. For example, Zhu et al. [32] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices. Huang et al. [33] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection.
Although these lightweight models reduce computational resources, their emphasis on minimizing parameters results in limited model expressiveness, hindering their ability to effectively capture intricate patterns and features in the input data. This leads to unsatisfactory segmentation results. This imbalance between performance and lightweight design makes them unreliable in practical applications.
III Methodology
III-A Network Architecture
As shown in Fig. 2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM). The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled. In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
Specifically, for a given dataset, the crack images are first input into the MFE module, which is divided into four stages. In each stage, the extracted feature maps undergo a Feature Enhancement stage. Then, the LDE module processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies. Subsequently, the SCFM module receives these sequences along with the four layers of feature maps extracted by the MFE. After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
Importantly, as shown in Fig. 3, using the original convolution operations results in high FLOPs and Params, especially in the SCFM module, reaching 73.36G and 9.71M, respectively. Replacing all convolutions with LRDS blocks significantly reduced FLOPs and Params, demonstrating the effectiveness of LRDS in model lightweighting.
Figure 3: Comparison of FLOPs and Params for each module using the original convolution and our proposed LRDS convolution block.
III-B Lightweight Multi-scale Feature Extractor
As shown in Fig. 2 and Fig. 4, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $C_{m}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $C_{d}$, producing the final output feature map.
Low-rank approximation decomposes a high-dimensional matrix into low-dimensional matrices, reducing computation complexity. In CNNs, it is applied to convolutional kernel responses, assuming these lie in a low-rank subspace.
Specifically, in the convolutional layer, we assume that its filter has a spatial size of $k$ and a number of input channels of $c$. In computing the convolutional response, the filter is applied to a sub-tensor of the input tensor of shape $k\times k\times c$. Suppose this sub-tensor is $t\in\mathbb{R}^{k^{2}c+1}$, where we append a bias term at the end. The response at a position in the output, $y\in\mathbb{R}^{e}$, can be expressed as follows:
$y=Wt$ (1)
where $e$ represents the number of filters and $W$ is a $e\times(k^{2}c+1)$ matrix. Assuming the vector $y$ lies in a low-rank subspace, we can write it as $y=U(y-y_{1})+y_{1}$, where $U$ is a $e\times e$ matrix of rank $e_{0}$, with $e_{0}<e$. $y_{1}$ is the mean vector of the response. Therefore, we can write the response as
$y=UWt+b$ (2)
where $b=y_{1}-Uy_{1}$ is the new bias. The rank of $U$ is $e_{0}$, and it can be decomposed into two $e\times e_{0}$ matrices $J$ and $K$ such that $U=JK^{\top}$. We use $W_{1}=K^{\top}W$ to denote a $e_{0}\times(k^{2}c+1)$ matrix, which corresponds to a set of $e_{0}$ filters. Thus, we can rewrite the response as follows:
$y=JW_{1}t+b$ (3)
The computational complexity of using Equation (3) is $O(e_{0}k^{2}c)+O(ee_{0})$, whereas approximating $y$ in the low-rank subspace using Equation (1) has a computational complexity of $O(ek^{2}c)$. Given that $O(ee_{0})\ll O(e_{0}k^{2}c)$, the complexity of using Equation (3) is $\frac{e_{0}}{e}$ times that of using Equation (1).
Besides low-rank approximation, which reduces convolutional layer parameters and computational load, depthwise separable convolutions [28] also significantly reduce computational resources by decoupling spatial and channel dimensions. Therefore, we embedded depthwise and pointwise convolution layers in the LRDS block. Additionally, to better preserve local contextual information during feature extraction while avoiding excessive computational overhead, we incorporated strip pooling [34] technology into each bottleneck of ResNet50.
Figure 4: Illustration of LRDS convolution block.
We constructed MFE by replacing all convolution operations in ResNet50 with LRDS convolution blocks.
III-C Long-range Dependency Extractor
Figure 5: Analysis of Computational Resources Required for Linear Operations. (a) shows the percentage of FLOPs and Params in the LDE module that use Original linear operations, with the inner circle showing the percentage that use FLOPs and the outer circle showing the percentage of Params. (b) Shows the comparison of FLOPs and Params for LDE before and after using LR Linear.
When extracting long-range dependencies of crack pixels, the deformable attention mechanism [35] adapts well to irregular structures. It calculates attention weights within a few sampled points, significantly reducing computational and memory overhead. Specifically, given an input $x\in\mathbb{R}^{C\times H\times W}$, let $q$ denote a query element with query feature $z_{q}$ and reference point coordinates $(p_{qx},p_{qy})$, the deformable attention is calculated as:
$\centering\begin{split}&\text{MSDeformAttn}(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg{[}\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg{]}\end{split}\@add@centering$ (4)
Where $h$ represents the attention head, $c$ represents the input feature layer, and $t$ represents the sampling key point. $\Delta p_{hcqt}$ and $A_{hcqt}$ represent the offset and attention weight of the $t$-th sampling point in the $c$-th feature map and $h$-th attention head, respectively. The normalization is such that $\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}=1$. Here, normalized coordinates $\hat{p}\in[0,1]$ are used to represent the position of a reference point in the normalized feature map $\phi_{c}(\hat{p}_{q})$ rescales the normalized coordinates onto the $c$-th layer feature map.
The structure of LDE is shown in Fig. 2. Specifically, before the enhanced feature maps enters the LDE encoder, it is flattened into a long pixel sequence, aiding the deformable attention mechanism in extracting pixel dependencies. As shown in Fig. 5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module. To reduce these, we designed a lightweight Linear layer (LR Linear) using low-rank approximation, as shown in Fig. 2. We split the original Linear operation into two parts: $f_{m}$ passes through the Reduce layer to get $f_{mid}$ with 4 channels, then through the Expand layer to produce $f_{out}$. Replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M, as shown in Fig. 5(b).
III-D Staircase Cascaded Fusion Module
Figure 6: Illustration of the PAF and SeFu.
The design of our SCFM structure is shown in Fig. 2. To enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we improved the Pixel Attention Guided Fusion (PAF) module, inspired by PIDNet [36] and the need for model lightweighting. The structure of PAF is shown in Fig. 6(a). We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively. Thus, the output can be expressed as:
$\sigma=\text{Sigmoid}(f_{c}(\vec{v_{c}})\cdot f_{d}(\vec{v_{d}}))$ (5)
where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. $\sigma$ indicates the probability that the two pixels belong to the same category. If $\sigma$ is high, the model will trust $\vec{v_{c}}$ more, indicating that its semantic features are more accurate, and vice versa. Thus, the output of PAF can be represented as::
$\text{Output}=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (6)
To improve inference speed and reduce computational resource usage, all convolution operations in this module utilize LRDS blocks. As seen in Fig. 3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively. Thus, lightweighting these operations is crucial. Replacing all convolution operations with LRDS blocks reduces FLOPs and Params by 92.87% and 97.22%, respectively, demonstrating the effectiveness of LRDS in network lightweighting. SCFM is divided into four stages, each with two fusion branches: the concat branch and the PAF branch. The concat operation retains complete information and diverse feature representations through direct concatenation, while the PAF module selectively fuses features through a pixel-guided attention mechanism, avoiding issues like buried feature information, information conflict, and redundancy. By perceiving the feature maps from the two branches, both local patterns of crack pixels and long-range dependencies can be comprehensively addressed.
To fully utilize the outputs of MFE and LDE and improve segmentation performance, SCFM incorporates the outputs from earlier stages and the associated MFE feature maps starting from the second stage. These are fused through the PAF module and processed by the Selective Feature Fusion (SeFu) module [7], which combines all PAF outputs before further operations. This approach effectively combines local semantic and long-range information, improving segmentation. The SeFu module structure is shown in Fig. 6(b).
III-E Loss Function
The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance. Previously, Binary Cross-Entropy (BCE) [37] was often used, but it may not effectively distinguish the minority crack pixels in the image. Combining BCE with Dice loss [38] addresses this issue. BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data.
Our loss $L$ can be calculated using Equation (7), which combines the above two loss functions for joint optimization.
$L=\alpha\cdot L_{BCE}+\beta\cdot L_{Dice}$ (7)
where $\alpha$ and $\beta$ are hyperparameters that determine the weights of the two losses. In our network, $\alpha$ is set to 0.75 and $\beta$ to 0.25.
IV Experiments
In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics. We then analyze results and model complexity, perform ablation experiments, and discuss current limitations and considerations.
IV-A Dataset Description
TABLE I: Comparison of Different Crack Datasets
No. Datasets Year Num(train/val/test) Resolution Percentage of cracked pixels Scenarios
1 CFD [39] 2016 118(76/11/31) 480 width and 320 height 1.60% bitumen
2 GAPS509 [40] 2017 509(307/100/102) 540 width and 640 height 1.19% bitumen
3 DeepCrack [21] 2019 537(300/100/247) 544 width and 384 height 3.54% bitumen, cement
4 Crack500 [22] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen
5 CrackMap [41] 2023 120(72/12/36) 256 width and 256 height 4.10% bitumen
6 AEL [22] 2019 58 768 width and 512 height 0.67% bitumen
7 CrackTree206 [42] 2012 206 800 width and 600 height 0.32% bitumen
8 LSCD [43] 2024 143 400 width and 400 height 3.18% bitumen, cement
9 CrackSC [44] 2023 197 320 width and 480 height 2.02% bitumen, cement
10 TUT (Ours) 2024 1408(987/139/282) 640 width and 640 height 3.16% bitumen, cement, bricks, runways, metal, blades, pipelines, tiles
Figure 7: Illustration of the images of each scenarios in the TUT dataset.
The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only. These datasets have clear backgrounds and minimal interference, which cannot comprehensively evaluate the performance of models. To tackle this problem, we created the TUT dataset to enable better segmentation across different scenarios. Detailed information and comparisons with other datasets are list in Table I.
The TUT dataset includes a substantial number of crack images captured with mobile phones, along with a smaller collection of images sourced from the internet. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex, noisy backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
As list in Table I, our collected TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Fig. 7 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.
As seen in Fig. 7, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties. Particularly in plastic runways, both coarse and fine cracks are present, testing the model’s generalization to different types of cracks. In images of tiles, metal materials, wind turbine blades, and underground pipelines, the crack shapes are relatively simple, but the backgrounds are very complex with various external interferences such as low light, irrelevant areas, and surface characteristics. This can easily result in false detections, thereby testing the model’s ability to accurately extract crack areas from complex environments.
IV-B Implementation Details
1) Experimental Settings: The CrackSCF network is built on the PyTorch v2.1.1 deep learning framework and trained on an Ubuntu 20.04 server, featuring two Intel(R) Xeon(R) Platinum 8336C CPUs and eight Nvidia GeForce 4090 GPUs, each with 24GB VRAM. For training, the AdamW optimizer is utilized with an initial learning rate set at 0.0005, weight decay of 0.0001, batch size of one, and a total of 60 training epochs. The learning rate is reduced to one-tenth after 30 epochs as part of a decay strategy. The best-performing model parameters on the validation set are selected for testing. We compared the CrackSCF network with other SOTA methods on six datasets: CFD [39], GAPS509 [40], DeepCrack [21], Crack500 [22], CrackMap [41], and ours TUT dataset.
2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation algorithms and the latest SOTA methods on different datasets. These methods include HED [45], UNet++ [46], Deeplabv3+ [47], AttuNet [48], RCF [49], RIND [50], UCTNet [6], SFIAN [7], CT-crackseg [8], DTrCNet [9], Crackmer [10] and Simcrack [11].
3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), Mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS). Mean Intersection over Union (mIoU) measures the average ratio of the intersection and union of the predicted results and the true results. The calculation formula for mIoU is:
$MIoU=\frac{1}{i+1}\sum_{a=0}^{i}\frac{p_{aa}}{\sum_{b=0}^{i}p_{ab}+\sum_{b=0}^{i}p_{ba}-p_{aa}}$ (8)
where $i$ is the number of classes, $a$ denotes the true value, $b$ denotes the predicted value, and $p_{ab}$ represents the number of pixels predicted as $a$ for $b$. Given that our crack segmentation task involves binary classification, we set $i=1$.
Optimal Dataset Scale (ODS) indicates the optimal segmentation performance metric obtained when a fixed threshold $t$ is selected over the entire dataset. It is defined as:
$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (9)
Optimal Image Scale (OIS) indicates the optimal segmentation performance metric obtained when an optimal threshold $t$ is selected for each image individually. It is defined as:
$OIS=\frac{1}{N}\sum_{a=1}^{N}\max_{t}\frac{2\cdot P_{t,a}\cdot R_{t,a}}{P_{t,a}+R_{t,a}}$ (10)
where $N$ represents the total number of images.
IV-C Comparison with the SOTA methods
TABLE II: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Results are highlighted with the best in red and the second-best in blue
Methods CFD GAPS509 Crack500
ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU
HED 0.6276 0.6494 0.5467 0.7374 0.6279 0.7240 0.4014 0.4509 0.4300 0.4624 0.4456 0.6282 0.6191 0.6537 0.6150 0.7281 0.6668 0.7140
UNet++ 0.6586 0.6836 0.5995 0.7608 0.6706 0.7432 0.6255 0.6484 0.6160 0.6684 0.6411 0.7300 0.6965 0.7274 0.6989 0.7805 0.7374 0.7605
DeeplabV3+ 0.6910 0.7084 0.6264 0.7730 0.6920 0.7709 0.6160 0.6425 0.6050 0.6669 0.6345 0.7378 0.6943 0.7211 0.6970 0.7569 0.7257 0.7746
AttuNet 0.6710 0.6870 0.6240 0.7220 0.6694 0.7576 0.5955 0.6174 0.5924 0.6519 0.6208 0.7247 0.6586 0.6890 0.6490 0.7608 0.7005 0.7543
RCF 0.6434 0.6624 0.5792 0.7348 0.6478 0.7343 0.4532 0.4817 0.4420 0.5428 0.4872 0.6472 0.6054 0.6361 0.6037 0.7136 0.6540 0.7068
RIND 0.6858 0.7021 0.6456 0.7534 0.6954 0.7592 0.5505 0.6023 0.4959 0.6494 0.5623 0.6909 0.6469 0.6483 0.6998 0.7245 0.7119 0.7381
UCTNet 0.6587 0.7006 0.6701 0.7205 0.6944 0.7479 0.5835 0.6188 0.5709 0.6752 0.6187 0.7107 0.7147 0.7265 0.6887 0.7843 0.7302 0.7674
SFIAN 0.7040 0.7159 0.6657 0.7512 0.7058 0.7684 0.6087 0.6320 0.6061 0.6451 0.6250 0.7198 0.6977 0.7348 0.6983 0.7742 0.7343 0.7604
CT-crackseg 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
DTrCNet 0.7079 0.7123 0.6774 0.7486 0.7112 0.7710 0.6211 0.6276 0.6277 0.6592 0.6431 0.7272 0.7012 0.7241 0.6527 0.8280 0.7357 0.7627
Crackmer 0.6751 0.6754 0.6370 0.7237 0.6776 0.7503 0.4593 0.4642 0.6144 0.4663 0.5302 0.6581 0.6933 0.7097 0.6985 0.7572 0.7267 0.7591
SimCrack 0.7152 0.7204 0.6883 0.7557 0.7204 0.7756 0.6408 0.6471 0.6171 0.6997 0.6558 0.7368 0.7127 0.7308 0.7093 0.7984 0.7516 0.7715
Ours 0.7335 0.7376 0.6967 0.7848 0.7383 0.7864 0.6492 0.6552 0.6459 0.6812 0.6631 0.7411 0.7226 0.7353 0.7213 0.7922 0.7551 0.7764
CrackMap DeepCrack TUT (Ours)
HED 0.6022 0.6240 0.5281 0.7063 0.6043 0.7026 0.7552 0.8237 0.7231 0.7678 0.7448 0.8007 0.6960 0.7254 0.7080 0.7431 0.7251 0.7683
UNet++ 0.7030 0.7222 0.6773 0.7209 0.6984 0.7631 0.8419 0.8893 0.8474 0.8533 0.8503 0.8621 0.7702 0.7948 0.7725 0.7974 0.7848 0.8160
DeeplabV3+ 0.6617 0.6793 0.5853 0.7532 0.6587 0.7460 0.7727 0.8354 0.7363 0.8323 0.7813 0.8565 0.7504 0.7777 0.7155 0.8231 0.7655 0.8177
AttuNet 0.6618 0.6773 0.6233 0.6970 0.6581 0.7454 0.8153 0.8379 0.8066 0.8069 0.8068 0.8578 0.7539 0.7795 0.7558 0.7945 0.7747 0.8204
RCF 0.5552 0.5683 0.4590 0.7164 0.5595 0.6758 0.7666 0.8376 0.7423 0.7611 0.7516 0.8110 0.6821 0.7138 0.7106 0.6877 0.6990 0.7598
RIND 0.6745 0.6943 0.6023 0.7586 0.6699 0.7425 0.8087 0.8267 0.7896 0.8920 0.8377 0.8391 0.7531 0.7891 0.7872 0.7665 0.7767 0.8051
UCTNet 0.6104 0.6195 0.5653 0.6748 0.6152 0.7079 0.8357 0.8579 0.8217 0.8857 0.8525 0.8564 0.7569 0.7670 0.8201 0.7667 0.7925 0.8110
SFIAN 0.7200 0.7465 0.6715 0.7668 0.7160 0.7748 0.8616 0.8928 0.8549 0.8692 0.8620 0.8776 0.7290 0.7513 0.7715 0.7367 0.7537 0.7896
CT-crackseg 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
DTrCNet 0.7328 0.7413 0.6912 0.7681 0.7276 0.7812 0.8473 0.8512 0.8905 0.8251 0.8566 0.8661 0.7987 0.8073 0.7972 0.8441 0.8202 0.8331
Crackmer 0.7395 0.7437 0.7229 0.7467 0.7346 0.7860 0.8712 0.8785 0.8946 0.8783 0.8864 0.8844 0.7429 0.7640 0.7501 0.7656 0.7578 0.7966
SimCrack 0.7559 0.7625 0.7380 0.7672 0.7523 0.7963 0.8570 0.8722 0.8984 0.8549 0.8761 0.8744 0.7984 0.8090 0.8051 0.8371 0.8208 0.8334
Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
Our proposed method, CrackSCF, was compared with 12 other methods across five public datasets and our own TUT dataset. Table II displays the comparison results for the five public datasets, while Fig. 8 illustrates the visual results of all methods on these datasets.
Figure 8: Visualization results of CrackSCF and other methods on public datasets. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
1) Results on the CFD dataset: The crack areas in the CFD dataset are very slender, testing the model’s sensitivity to fine-grained crack pixels. As list in Table II, our proposed CrackSCF attained the best results across all metrics for the CFD dataset. Specifically, on ODS, OIS, P, R, F1, and mIoU, it outperformed the second-best methods by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively. Notably, SimCrack [11] achieved the second-best performance on all metrics except for P and R, only behind our method. However, as seen in Fig. 8, its actual segmentation performance is not satisfactory, while our method achieves high-quality segmentation results on various fine-grained crack pixels. Other methods, such as UCTNet [6] and DTrCNet [9], tend to misdetect background noise or miss key detail areas of crack pixels.
2) Results on the GAPS509 dataset: Unlike the CFD dataset, the GAPS509 dataset contains not only finer cracks but also some irrelevant background interference, testing the model’s ability to extract crack areas under high interference conditions. As list in Table II, our method achieved the highest performance on all metrics except R. Although SimCrack [11] achieved 0.6997 on the R metric, leading our method by about 2.72%, its P metric was 0.6171, far lower than our method. The highest F1 score obtained by our method indicates a good balance between P and R. Specifically, our method outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.31%, 1.25%, 2.90%, 1.11%, and 0.45%, respectively. Additionally, as seen in Fig. 8, our method almost perfectly avoids irrelevant background noise, while most other methods falsely detect irrelevant backgrounds and perform worse in segmenting crack pixel areas compared to our method.
3) Results on the DeepCrack dataset: The cracks in the DeepCrack dataset are often relatively wide, testing the model’s ability to extract a large range of crack pixels. As list in Table II, our method achieved the highest results on all metrics for the DeepCrack dataset. Specifically, our method outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85%, respectively. Notably, CT-crackseg also achieved good results on this dataset, but as seen in Fig. 8, its segmentation performance on detailed cracks is not as good as our method, indicating that our method can effectively handle crack details.
4) Results on the Crack500 dataset: Similar to the DeepCrack dataset, the Crack500 dataset also has relatively wide crack areas, but its structure is more complex, posing higher requirements on model performance. As list in Table II, our method achieved the best results on five metrics for the Crack500 dataset, except for R. It outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%, respectively. Although DTrCNet [9] leads on the R metric, our method surpasses it on the P metric by 10.51%. The highest F1 score achieved by our method indicates a good balance between P and R. Additionally, as seen in Fig. 8, our method does not produce false detections in irrelevant background areas, whereas other methods either produce false detections or lack detailed segmentation in crack areas.
5) Results on the CrackMap dataset: Similar to the CFD dataset, the cracks in the CrackMap dataset are also relatively slender. As list in Table II, our method leads all other methods on all metrics, outperforming the second-best methods on ODS, OIS, P, R, F1, and mIoU by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively. Notably, SimCrack [11] achieved the second-best results on this dataset, but as seen in Fig. 8, SimCrack’s segmentation performance on the crack tips is not satisfactory, whereas our method effectively identifies these pixels.
Figure 9: Visualization results of CrackSCF and other methods on TUT dataset. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
6) Results on the TUT dataset: Fig. 9 shows the visual results of all methods on the TUT dataset. We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes. As list in Table II, our method achieved excellent results on the TUT dataset, leading in all metrics. Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively. SimCrack [11] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig. 9, its segmentation performance on some key details has room for improvement compared to our method.
Furthermore, most methods fail to suppress irrelevant background noise. In images of metal materials and turbine blades, our method effectively removes noise and accurately segments crack regions. In pipeline scenarios, our method can filter out timestamp watermarks, thanks to SCFM, which enables the model to focus on irregular, slender crack structures rather than irrelevant watermarks. Our method demonstrates excellent robustness and versatility, resulting in high-quality crack segmentation maps and adapting well to different scenarios.
IV-D Complexity Analysis
To assess our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS). FLOPs quantify computational load, reflecting demand during inference. Params measure storage requirements, reflecting trainable and stored parameters. FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
Table III presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [9]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines minimal parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
TABLE III: Comparison of the complexity of CrackSCF with other methods. Results are highlighted with the best in red and the second-best in blue
Methods FLOPs ↓ Params ↓ FPS ↑
HED 80.35G 14.72M 33
RCF 115.54G 15.52M 30
RIND 695.77G 59.39M 11
UCTNet 171.77G 66.22M 12
UNet++ 139.61G 9.16M 22
DeeplabV3+ 88.53G 59.34M 29
AttuNet 541.34G 57.16M 24
SFIAN 52.36G 13.63M 35
CT-crackseg 39.47G 22.88M 28
DTrCNet 123.20G 63.45M 47
SimCrack 70.66G 29.58M 29
Crackmer 14.94G 5.90M 31
Ours 9.26G 4.79M 36
TABLE IV: Comparison of CrackSCF using LRDS blocks with networks using other convolutions. Results are highlighted with the best in red and the second-best in blue
Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
CrackSCF_Original 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
CrackSCF_DS 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
CrackSCF_LR 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
CrackSCF_LRDS 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE V: Comparison of CrackSCF Networks using SCFM, Concat and PAF. Results are highlighted with the best in red and the second-best in blue
Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
CrackSCF_Concat 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359 7.21G 4.71M 42
CrackSCF_PAF 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427 8.24G 4.72M 39
CrackSCF_SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE VI: Comparison of results of CrackSCF networks trained with different loss ratios
Dice : BCE ODS OIS P R F1 mIoU
1 : 5 0.8129 0.8203 0.8217 0.8458 0.8336 0.8424
1 : 4 0.8049 0.8145 0.8116 0.8365 0.8239 0.8368
1 : 3 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
1 : 2 0.8146 0.8211 0.8227 0.8435 0.8330 0.8438
1 : 1 0.8136 0.8204 0.8231 0.8436 0.8332 0.8431
Only BCE 0.8003 0.8056 0.8349 0.8004 0.8173 0.8332
Only Dice 0.8107 0.8165 0.8302 0.8265 0.8283 0.8409
IV-E Ablation Analysis
To evaluate the advantages of our LRDS block compared to other convolution operations, we performed ablation experiments on the TUT dataset, as list in Table IV. On six metrics, our method slightly outperforms those using only depthwise separable or low-rank approximation convolutions. Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the low-rank approximation block, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that our convolution block effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, we performed ablation experiments on the TUT dataset, as list in Table V. Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics. Using only the PAF branch results in a 1.06% drop in the F1 score. While SCFM slightly increases parameter count, computational load, and inference speed, it significantly improves segmentation performance. These results indicate that our module effectively balances performance and efficiency.
Table VI shows the results of ablation experiments on different weight proportions of the two sub-losses in the loss function. As shown, a Dice loss to BCE loss ratio of 1:3 yields the highest values in five metrics: ODS, OIS, R, F1, and mIoU. Although the P metric is 0.80% lower than using only BCE loss, the F1 metric improves by 2.58%, indicating a better balance between P and R. Therefore, we use this ratio during training. In summary, a Dice loss to BCE loss ratio of 1:3 not only enhances pixel-level performance but also improves the global segmentation of crack areas, achieving superior segmentation performance.
IV-F Limitations and Discussions
Through these experiments, we found that SCFM effectively perceives and fuses local details and long-range pixel dependencies. Our network excels in various metrics on public datasets with simple scenes and performs exceptionally on the complex TUT dataset. The lightweight LRDS block minimizes parameters and computational load, facilitating deployment on resource-constrained devices like drones and smartphones. However, our network has several limitations, as follows:
  1. 1.
    Although our method generally achieves good crack segmentation with effective background noise suppression, as shown in the Fig. 10, our model occasionally misdetects noise regions that resemble cracks in some images. It also fails to detect watermark occlusions with colors similar to the background. Therefore, it is essential to enhance modules such as MFE and LDE to improve the model’s capability to perceive and extract features from critical noise and occlusion areas, thereby further increasing its robustness and stability in complex scenarios.
  2. 2.
    Despite our method achieves the minimum parameter count and computational load, there is still room for improvement in inference speed. For example, our network achieves a processing speed of 36 frames per second, but the DTrCNet [9], which focuses more on inference speed, attains a faster speed of 47 frames per second. Therefore, it is necessary to further simplify convolution operations or the network’s complexity to achieve faster inference speeds while maintaining a good balance between segmentation performance and model lightweighting.
Figure 10: Examples of CrackSCF misdetections and misses. The critical regions are marked with blue boxes.
V Conclusion
This paper proposes a staircase cascaded fusion network (CrackSCF) for pixel-level crack detection. We developed SCFM to fuse the local patterns processed by MFE and the pixel dependencies processed by LDE. Additionally, we introduced the lightweight convolution block LRDS to replace all convolution operations in the network, effectively reducing the substantial parameter count and computational load required for training and inference. Furthermore, to comprehensively evaluate our model’s robustness and stability across various complex scenarios, we collected the TUT dataset, which includes eight types of crack scenes. Experimental results show that our method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation. The computational load and parameter count are reduced to 9.26G and 4.79M, respectively, with an inference speed of up to 36 frames per second.
In future work, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results. Additionally, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications.

### v3

@@
 Abstract
-Detecting cracks with pixel-level precision for key structures is a significant challenge, as existing methods struggle to effectively integrate local textures and pixel dependencies of cracks.
-Furthermore, these methods often possess numerous parameters and substantial computational requirements, complicating deployment on edge control devices.
-In this paper, we propose a staircase cascaded fusion crack segmentation network (CrackSCF) that generates high-quality crack segmentation maps using minimal computational resources.
-We constructed a staircase cascaded fusion module that effectively captures local patterns of cracks and long-range dependencies of pixels, and it can suppress background noise well.
-To reduce the computational resources required by the model, we introduced a lightweight convolution block, which replaces all convolution operations in the network, significantly reducing the required computation and parameters without affecting the network’s performance.
-To evaluate our method, we created a challenging benchmark dataset called TUT and conducted experiments on this dataset and five other public datasets.
-The experimental results indicate that our method offers significant advantages over existing methods, especially in handling background noise interference and detailed crack segmentation.
-The F1 and mIoU scores on the TUT dataset are 0.8382 and 0.8473, respectively, achieving state-of-the-art (SOTA) performance while requiring the least computational resources.
+Detecting cracks with pixel-level precision for key structures is a significant challenge, existing methods struggle to integrate local textures and pixel dependencies of cracks.
+Furthermore, these methods possess numerous parameters and substantial computational requirements, complicating deployment on edge devices.
+In this paper, we propose the Staircase Cascaded Fusion Crack Segmentation Network (CrackSCF), which generates high-quality crack segmentation maps while reducing computational overhead.
+We design a lightweight convolutional block that substitutes all convolution operations, reducing the model’s computational demands while maintaining an effective capture of local details.
+Additionally, we introduce a lightweight long-range dependency extractor to better capture the long-range dependencies.
+Furthermore, we develop a staircase cascaded fusion module, which seamlessly integrates local patterns and long-range dependencies, resulting in high-quality segmentation maps.
+To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets.
+The results show that our method outperforms existing methods, particularly in handling background noise and achieving detailed segmentation.
+The F1 and mIoU scores on the TUT dataset are 0.8382 and 0.8473, respectively, demonstrating state-of-the-art (SOTA) performance with low computational resources.
 The code and dataset is available at https://github.com/Karl1109/CrackSCF.
 Index Terms:
-Structural cracks segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
+Structural crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
+publicationid: pubid: $0$–$00$ 2025 IEEE
 I Introduction
 Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues.
-Thus, automated detection on edge control devices is crucial for ensuring production safety [1, 2].
+Thus, automated detection on edge control devices is crucial for ensuring production safety [1].
 Early automated detection methods used traditional image processing techniques for crack extraction [2, 3, 4].
@@
 Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
-The success of VIT [14] and Swin Transformer [15] has demonstrated the advantages of Transformer [16], with their strong sequence processing capabilities, in handling visual tasks.
+The success of Vision Transformer (ViT) [13] and Swin Transformer [14] has demonstrated the advantages of Transformer [15], with their strong sequence processing capabilities, in handling visual tasks [16, 17].
 Unlike CNNs, transformers can capture dependencies between any pixels when processing images [18, 19], making them more advantageous in handling complex topologies.
-To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [19], MFAFNet [20], and Crackmer [9], which integrate local texture information and long-range pixel dependencies through U-shaped connections or residual connections.
+To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [20], MFAFNet [21], and Crackmer [8], which integrate local and long-range information through U-shaped connections or residual connections.
 Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations, as follows:
@@
 1(c), networks like SFIAN [6], Crackmer [9], and CT-crackseg[7] have very high computational FLOPs and parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
-To address these issues, we propose a staircase cascaded fusion network that can generate high-quality segmentation maps with minimal computational resources while effectively suppressing background noise.
-Specifically, to better exploit the complementarity and correlation of feature maps without missing critical detail information, we introduce a staircase cascaded fusion module.
-This module utilizes the pixel attention mechanism, which finely captures local and long-range information, and the channel concatenation mechanism, which progressively integrates features from adjacent layers, to generate crack segmentation maps that comprehensively combine local texture information and pixel dependencies.
-Furthermore, to decrease the parameter count and computational burden, we introduce a lightweight convolution block called LRDS.
-This block initially increases and then reduces the channel count, and separates the spatial and channel aspects, thereby minimizing computational costs while maintaining performance.
-To assess the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images.
+To address these issues, we propose a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise.
+Specifically, to reduce the number of parameters and computational burden, we introduce a lightweight convolution block called LRDS.
+This module utilizes the principle of low-rank approximation to effectively reduce the computational resources required by convolutional layers, and captures spatial structures efficiently in the low-rank space through depthwise convolution, maintaining good feature perception while requiring very low computational resources.
+We also propose a lightweight Long-range Dependency Extractor (LDE), which uses a lightweight deformable attention mechanism to effectively capture irregular long-range dependencies between crack pixels with very low computational demands.
+Additionally, to better leverage the complementarity and correlation of feature maps while ensuring that critical detail information is not missed, we introduce the Staircase Cascaded Fusion Module (SCFM).
+This module generates feature maps layer by layer and effectively fuses them using a pixel attention mechanism that can finely capture local and long-range information and a channel concatenation mechanism, comprehensively combining local texture information and pixel dependencies.
+To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images.
 Existing public datasets like DeepCrack [22] and Crack500 [23] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance.
 The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines.
-With such a diverse set of image scenarios, this dataset effectively assesses the network’s generalization and robustness.
+With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness.
 Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
@@
     We proposed a novel CrackSCF network for structural crack segmentation.
-By introducing a staircase cascaded fusion module and a lightweight convolution block, this network effectively distinguishes between background and crack pixels with minimal computational resources, resulting in high-quality segmentation maps.
+By introducing a lightweight convolution block, a lightweight long-range dependency extractor, and a staircase cascaded fusion module, this network effectively distinguishes between background and crack pixels with low computational resources, resulting in high-quality segmentation maps.
   2. 2.
@@
 II Related Work
+Figure 2: General Architecture Diagram of CrackSCF Network.
+Crack images are first input into the MFE, generating four feature maps.
+After enhancement, five feature maps with unified channel numbers are obtained.
+The LDE processes these maps to acquire pixel sequences rich in long-range correlations.
+The SCFM then processes these sequences and the four layers of feature maps from the MFE.
+Through four stages, resulting in the segmentation output.
 II-A Crack Segmentation Networks
@@
 [25] developed a crack detection network for masonry structures that combines thermal imaging and visible light image fusion with semantic segmentation, laying the foundation for the development of cross-modal methods for structural crack monitoring.
-Although the aforementioned methods have demonstrated the advantages of CNNs in crack segmentation tasks, CNNs are essentially local operations and cannot effectively perceive a small number of crack pixels amidst a large background.
+Although the aforementioned methods have demonstrated the advantages of CNNs in crack segmentation tasks, CNNs are essentially local operations and cannot effectively perceive a small number of pixels amidst a large background.
 Therefore, these methods often result in discontinuous crack pixels and are highly susceptible to background noise interference.
-Figure 2: General Architecture Diagram of CrackSCF Network.
-Crack images are first input into the MFE, generating four feature maps.
-After enhancement, five feature maps with unified channel numbers are obtained.
-The LDE module processes these maps to acquire pixel sequences rich in long-range correlations.
-The SCFM module then processes these sequences and the four layers of feature maps from the MFE.
-Through four stages, resulting in the segmentation output.
-Thanks to the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes.
+Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes.
 Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation.
 Chen et al.
-[25] proposed TransUNet, a novel network architecture that effectively integrates long-range information and local features through a U-Net-like structure.
+[26] proposed TransUNet, a novel network that effectively integrates long-range information and local features through a U-Net-like structure.
 Qi et al.
-[26] developed an end-to-end model based on Vision Transformer (ViT) and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders.
+[27] developed an end-to-end model based on ViT and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders.
 Feng et al.
@@
 However, these devices have limited computing resources, making it challenging to use models with high parameter counts.
-Thus, compact model design for lightweight modifications in size and speed is crucial.
+Thus, compact model design for lightweight modifications in size and speed is crucial [29, 30].
 For example, Howard et al.
@@
 Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures.
-Zhang et al.
-[30] proposed a low-rank technique for neural networks that employs constrained optimization to to determine the optimal low-rank approximation.
 Swaminathan et al.
 [33] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
-Recently, many researchers have combined model lightweighting techniques with crack segmentation models to enable real-time crack detection on edge devices.
+In recent years, there has been a growing trend among researchers to integrate model lightweighting techniques with crack segmentation models, aiming to facilitate real-time crack detection on edge computing devices.
 For example, Zhu et al.
@@
 As shown in Fig.
-2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM).
+2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a lightweight Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM).
 The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled.
@@
 As shown in Fig.
-2 and Fig.
 4, our designed LRDS block adopts a bottleneck-like structure.
@@
 Given that $O(ee_{0})\ll O(e_{0}k^{2}c)$, the complexity of using Equation (3) is $\frac{e_{0}}{e}$ times that of using Equation (1).
-Besides low-rank approximation, which reduces convolutional layer parameters and computational load, depthwise separable convolutions [28] also significantly reduce computational resources by decoupling spatial and channel dimensions.
-Therefore, we embedded depthwise and pointwise convolution layers in the LRDS block.
-Additionally, to better preserve local contextual information during feature extraction while avoiding excessive computational overhead, we incorporated strip pooling [34] technology into each bottleneck of ResNet50.
+Low-rank approximation theory greatly reduces the convolutional layer parameters and computational load.
+In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space.
+We constructed MFE by replacing all convolution operations in ResNet50 with LRDS convolution blocks.
 Figure 4: Illustration of LRDS convolution block.
-We constructed MFE by replacing all convolution operations in ResNet50 with LRDS convolution blocks.
-III-C Long-range Dependency Extractor
+After the number of channels of the input features is reduced, their local and spatial features are efficiently extracted by depth-wise convolution and point-wise convolution, and finally the number of channels is restored.
+III-C Lightweight Long-range Dependency Extractor
 Figure 5: Analysis of Computational Resources Required for Linear Operations.
@@
 (b) Shows the comparison of FLOPs and Params for LDE before and after using LR Linear.
-When extracting long-range dependencies of crack pixels, the deformable attention mechanism [35] adapts well to irregular structures.
-It calculates attention weights within a few sampled points, significantly reducing computational and memory overhead.
+When extracting long-range dependencies of crack pixels, the deformable attention mechanism adapts well to irregular structures.
 Specifically, given an input $x\in\mathbb{R}^{C\times H\times W}$, let $q$ denote a query element with query feature $z_{q}$ and reference point coordinates $(p_{qx},p_{qy})$, the deformable attention is calculated as:
@@
 Here, normalized coordinates $\hat{p}\in[0,1]$ are used to represent the position of a reference point in the normalized feature map $\phi_{c}(\hat{p}_{q})$ rescales the normalized coordinates onto the $c$-th layer feature map.
+As shown in Fig.
+5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module.
+In order to reduce the computational resources required by the LDE, we propose a linear transformation method (LR Linear) based on low-rank approximation.
+In traditional linear layers, the weight matrix connecting the input and output is typically a high-rank matrix, representing a complete connection between input and output channels.
+By applying low-rank approximation, we decompose the weight matrix into two smaller matrices, reducing the number of parameters.
+Specifically, the weight matrix $E\in\mathbb{R}^{r_{\text{out}}\times r_{\text{in}}}$, where $r_{\text{out}}$ is the output feature dimension and $r_{\text{in}}$ is the input feature dimension, is approximated by two smaller matrices $A\in\mathbb{R}^{r_{\text{out}}\times r}$ and $B\in\mathbb{R}^{r_{\text{in}}\times r}$, where $r$ is a small rank, much smaller than $r_{\text{out}}$ and $r_{\text{in}}$.
+This decomposition breaks the original full connection into two smaller linear transformations:
+$E\approx E^{\prime}=AB^{T}$ (5)
+we perform this transformation through two linear layers.
+First, we apply a linear transformation $B^{T}$ to the input $x\in\mathbb{R}^{r_{\text{in}}}$, resulting in an intermediate representation $z\in\mathbb{R}^{r}$:
+$z=B^{T}x$ (6)
+Then, we apply another linear transformation $A$ to $z$, resulting in the output $y\in\mathbb{R}^{r_{\text{out}}}$:
+$y=Az$ (7)
+As $r\ll r_{\text{out}},r_{\text{in}}$, the low-rank approximation reduces the original matrix to two smaller matrices, significantly reducing the computational and storage requirements.
+This leads to a reduction in both the computational complexity and the number of parameters, thus improving training and inference efficiency.
+As shown in Fig.
+5(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
 The structure of LDE is shown in Fig.
 2.
-Specifically, before the enhanced feature maps enters the LDE encoder, it is flattened into a long pixel sequence, aiding the deformable attention mechanism in extracting pixel dependencies.
-As shown in Fig.
-5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module.
-To reduce these, we designed a lightweight Linear layer (LR Linear) using low-rank approximation, as shown in Fig.
-2.
-We split the original Linear operation into two parts: $f_{m}$ passes through the Reduce layer to get $f_{mid}$ with 4 channels, then through the Expand layer to produce $f_{out}$.
-Replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M, as shown in Fig.
-5(b).
+LDE facilitates long-range dependency modeling through the interaction of multi-scale features between the encoder and decoder.
+Initially, the encoder inputs a multi-scale feature set $\{X_{c}\}_{c=1}^{C}$ (where $C$ represents the number of feature levels).
+After undergoing layer flattening and positional encoding enhancement, the encoder extracts the feature sequence $S\in\mathbb{R}^{B\times L\times d}$ (where $L$ denotes the sequence length).
+The encoder then combines LR Linear with lightweight deformable attention mechanisms to capture long-range dependencies:
+$Z=\text{MSDeformAttn}(S,P,S)$ (8)
+where $P\in\mathbb{R}^{B\times L\times 2}$ represent the normalized reference coordinates.
+Subsequently, $Z$ is processed through a feedforward network (FFN) that combines LR Linear to refine the features:
+$M=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Z)))$ (9)
+where the output memory matrix $M\in\mathbb{R}^{B\times L\times d}$ is used as the input to the decoder, encoding multi-scale local dependencies.
+The decoder processes the input with a learnable query vector $Q\in\mathbb{R}^{N\times d}$ and the initial reference point $H=\text{Sigmoid}(TQ)$, where $T\in\mathbb{R}^{d\times 2}$ is the projection matrix.
+First, the reference points $H\in\mathbb{R}^{N\times 2}$ are mapped to the encoded memory $M$ to aggregate features:
+$Y_{0}=\text{MSDeformAttn}(Q,H,M)$ (10)
+Then, it passes through the FFN to obtain the crack pixel sequence enriched with irregular long-range dependencies:
+$Y=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Y_{0})))$ (11)
+where Y is the output sequence of the LDE enriched with irregular long-range dependencies between pixels.
 III-D Staircase Cascaded Fusion Module
-Figure 6: Illustration of the PAF and SeFu.
+Figure 6: Illustration of the PAF. It effectively reduces required computing resources by embedding LRDS.
 The design of our SCFM structure is shown in Fig.
 2.
-To enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we improved the Pixel Attention Guided Fusion (PAF) module, inspired by PIDNet [36] and the need for model lightweighting.
+As a segmentation head, to enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we proposed the lightweight Pixel Attention Guided Fusion (PAF) module.
 The structure of PAF is shown in Fig.
-6(a).
+6.
 We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively.
-Thus, the output can be expressed as:
+The output can be expressed as:
 $\sigma=\text{Sigmoid}(f_{c}(\vec{v_{c}})\cdot f_{d}(\vec{v_{d}}))$ (12)
-where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively.
-$\sigma$ indicates the probability that the two pixels belong to the same category.
+where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. Thus, the output of PAF can be represented as:
+$\text{Output}=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (13)
+where $\sigma$ indicates the probability that the two pixels belong to the same category.
 If $\sigma$ is high, the model will trust $\vec{v_{c}}$ more, indicating that its semantic features are more accurate, and vice versa.
-Thus, the output of PAF can be represented as::
-$\text{Output}=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (6)
-To improve inference speed and reduce computational resource usage, all convolution operations in this module utilize LRDS blocks.
+To reduce computational resource usage, all convolution operations in this module utilize LRDS blocks.
 As seen in Fig.
@@
 The concat operation retains complete information and diverse feature representations through direct concatenation, while the PAF module selectively fuses features through a pixel-guided attention mechanism, avoiding issues like buried feature information, information conflict, and redundancy.
+The processes of SCFM can be given by the following equation:
+$\begin{split}&F_{i}^{\text{O}}=C_{\text{i}}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\right)\odot P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\end{split}$ (14)
+Where $F_{i}^{\text{O}}$ represents the output features of the $i$-th SCFM layer ($i$ ranges from 0 to 3), $F_{i}^{\text{MFE}}$ represents the $i$-th layer MFE features, $C_{\text{i}}$ is an operation chain: Concat followed by LRDS, ReLU, and GN, $P_{i}$ is a PAF operation, $\mu(\cdot)$ is the sigmoid function generating spatial attention weights, and $\odot$ represents the Hadamard product.
+Specifically, when $i=0$, $F_{i-1}^{\text{O}}$ is defined as the output of LDE, i.e., $F_{-1}^{\text{O}}=\text{LDE}_{\text{output}}$, ensuring an initial output feature is provided at the beginning of the computation.
 By perceiving the feature maps from the two branches, both local patterns of crack pixels and long-range dependencies can be comprehensively addressed.
-To fully utilize the outputs of MFE and LDE and improve segmentation performance, SCFM incorporates the outputs from earlier stages and the associated MFE feature maps starting from the second stage.
-These are fused through the PAF module and processed by the Selective Feature Fusion (SeFu) module [7], which combines all PAF outputs before further operations.
-This approach effectively combines local semantic and long-range information, improving segmentation.
-The SeFu module structure is shown in Fig.
-6(b).
+To fully utilize the outputs of MFE and LDE and improve segmentation performance, SCFM utilized the property that PAF can process multiple feature maps simultaneously by merging the outputs of all the preceding phases and the associated MFE feature maps from the second phase onwards.
+The processing enables SCFM to more comprehensively dissect and process the local patterns and long-range dependencies in the different stages of the feature map to achieve a finer and more continuous segmentation map.
+Furthermore, the fine-grained features extracted with PAF can be perceived more finely through the dynamic weighting mechanism to achieve a clearer segmentation map.
 III-E Loss Function
@@
 Our loss $L$ can be calculated using Equation (15), which combines the above two loss functions for joint optimization.
-$L=\alpha\cdot L_{BCE}+\beta\cdot L_{Dice}$ (7)
-where $\alpha$ and $\beta$ are hyperparameters that determine the weights of the two losses. In our network, $\alpha$ is set to 0.75 and $\beta$ to 0.25.
+$L=0.75\cdot L_{B}+0.25\cdot L_{D}$ (15)
+where $\mathit{L_{B}}$ represents the BCE loss function and $\mathit{L_{D}}$ represents the Dice loss function.
 IV Experiments
 In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics.
-We then analyze results and model complexity, perform ablation experiments, and discuss current limitations and considerations.
+We then analyze results and model complexity, perform ablation experiments.
 IV-A Dataset Description
@@
 3 DeepCrack [22] 2019 537(300/100/247) 544 width and 384 height 3.54% bitumen, cement
-4 Crack500 [22] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen
+4 Crack500 [23] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen, cement
 5 CrackMap [40] 2023 120(72/12/36) 256 width and 256 height 4.10% bitumen
@@
 Detailed information and comparisons with other datasets are list in Table I.
-The TUT dataset includes a substantial number of crack images captured with mobile phones, along with a smaller collection of images sourced from the internet.
+The TUT dataset comprises 1270 crack images captured with our mobile phones, along with a small number of 138 images sourced from the Internet.
+TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances.
 To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images.
@@
 This evaluates the model’s capability to segment cracks in challenging conditions.
-As list in Table I, our collected TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios.
+As list in Table I, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios.
 Fig.
@@
 We compared the CrackSCF network with other SOTA methods on six datasets: CFD [38], GAPS509 [39], DeepCrack [22], Crack500 [23], CrackMap [40], and ours TUT dataset.
-2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation algorithms and the latest SOTA methods on different datasets.
+2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation methods and the latest SOTA methods on different datasets.
 These methods include HED [44], UNet++ [45], Deeplabv3+ [46], AttuNet [47], RCF [48], RIND [49], UCTNet [5], SFIAN [6], CT-crackseg [7], DTrCNet [8], Crackmer [9] and Simcrack [10].
-3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), Mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
-Mean Intersection over Union (mIoU) measures the average ratio of the intersection and union of the predicted results and the true results.
-The calculation formula for mIoU is:
-$MIoU=\frac{1}{i+1}\sum_{a=0}^{i}\frac{p_{aa}}{\sum_{b=0}^{i}p_{ab}+\sum_{b=0}^{i}p_{ba}-p_{aa}}$ (8)
-where $i$ is the number of classes, $a$ denotes the true value, $b$ denotes the predicted value, and $p_{ab}$ represents the number of pixels predicted as $a$ for $b$.
-Given that our crack segmentation task involves binary classification, we set $i=1$.
-Optimal Dataset Scale (ODS) indicates the optimal segmentation performance metric obtained when a fixed threshold $t$ is selected over the entire dataset. It is defined as:
-$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (9)
-Optimal Image Scale (OIS) indicates the optimal segmentation performance metric obtained when an optimal threshold $t$ is selected for each image individually. It is defined as:
-$OIS=\frac{1}{N}\sum_{a=1}^{N}\max_{t}\frac{2\cdot P_{t,a}\cdot R_{t,a}}{P_{t,a}+R_{t,a}}$ (10)
-where $N$ represents the total number of images.
-IV-C Comparison with the SOTA methods
-TABLE II: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Results are highlighted with the best in red and the second-best in blue
+TABLE II: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
 Methods CFD GAPS509 Crack500
@@
 Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
+3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), Mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
+Mean Intersection over Union (mIoU) measures the average ratio of the intersection and union of the predicted results and the true results.
+The calculation formula for mIoU is:
+$MIoU=\frac{1}{i+1}\sum_{a=0}^{i}\frac{p_{aa}}{\sum_{b=0}^{i}p_{ab}+\sum_{b=0}^{i}p_{ba}-p_{aa}}$ (16)
+where $i$ is the number of classes, $a$ denotes the true value, $b$ denotes the predicted value, and $p_{ab}$ represents the number of pixels predicted as $a$ for $b$.
+Given that our crack segmentation task involves binary classification, we set $i=1$.
+Optimal Dataset Scale (ODS) indicates the optimal segmentation performance metric obtained when a fixed threshold $t$ is selected over the entire dataset. It is defined as:
+$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (17)
+Optimal Image Scale (OIS) indicates the optimal segmentation performance metric obtained when an optimal threshold $t$ is selected for each image individually. It is defined as:
+$OIS=\frac{1}{N}\sum_{a=1}^{N}\max_{t}\frac{2\cdot P_{t,a}\cdot R_{t,a}}{P_{t,a}+R_{t,a}}$ (18)
+where $N$ represents the total number of images.
+IV-C Comparison with the SOTA methods
 Our proposed method, CrackSCF, was compared with 12 other methods across five public datasets and our own TUT dataset.
@@
 8, its segmentation performance on detailed cracks is not as good as our method, indicating that our method can effectively handle crack details.
-4) Results on the Crack500 dataset: Similar to the DeepCrack dataset, the Crack500 dataset also has relatively wide crack areas, but its structure is more complex, posing higher requirements on model performance.
+4) Results on the Crack500 dataset: Similar to the DeepCrack, the Crack500 dataset also has relatively wide crack areas, but its structure is complex, posing higher requirements on model performance.
 As list in Table II, our method achieved the best results on five metrics for the Crack500 dataset, except for R.
@@
 Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
-6) Results on the TUT dataset: Fig.
-9 shows the visual results of all methods on the TUT dataset.
-We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes.
-As list in Table II, our method achieved excellent results on the TUT dataset, leading in all metrics.
-Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively.
-SimCrack [11] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig.
-9, its segmentation performance on some key details has room for improvement compared to our method.
-Furthermore, most methods fail to suppress irrelevant background noise.
-In images of metal materials and turbine blades, our method effectively removes noise and accurately segments crack regions.
-In pipeline scenarios, our method can filter out timestamp watermarks, thanks to SCFM, which enables the model to focus on irregular, slender crack structures rather than irrelevant watermarks.
-Our method demonstrates excellent robustness and versatility, resulting in high-quality crack segmentation maps and adapting well to different scenarios.
-IV-D Complexity Analysis
-To assess our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS).
-FLOPs quantify computational load, reflecting demand during inference.
-Params measure storage requirements, reflecting trainable and stored parameters.
-FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
-Table III presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS.
-Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively.
-This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices.
-In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [9].
-This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications.
-Overall, our model combines minimal parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
-TABLE III: Comparison of the complexity of CrackSCF with other methods. Results are highlighted with the best in red and the second-best in blue
+TABLE III: Comparison of the complexity of CrackSCF with other methods. Best results are bolded and second best results are underlined
 Methods FLOPs ↓ Params ↓ FPS ↑
@@
 Ours 9.26G 4.79M 36
-TABLE IV: Comparison of CrackSCF using LRDS blocks with networks using other convolutions. Results are highlighted with the best in red and the second-best in blue
+6) Results on the TUT dataset: Fig.
+9 shows the visual results of all methods on the TUT dataset.
+We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes.
+As list in Table II, our method achieved excellent results on the TUT dataset, leading in all metrics.
+Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively.
+SimCrack [10] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig.
+9, its segmentation performance on some key details has room for improvement compared to our method.
+Furthermore, most methods fail to suppress irrelevant background noise.
+In images of metal materials and turbine blades, our method effectively removes noise and accurately segments crack regions.
+In pipeline scenarios, our method can filter out timestamp watermarks, thanks to SCFM, which enables the model to focus on irregular, slender crack structures rather than irrelevant watermarks.
+Our method demonstrates excellent robustness and versatility, resulting in high-quality crack segmentation maps and adapting well to different scenarios.
+TABLE IV: Comparison of CrackSCF using LRDS blocks with networks using other convolutions. Best results are bolded and second best results are underlined
 Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
-CrackSCF_Original 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
-CrackSCF_DS 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
-CrackSCF_LR 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
-CrackSCF_LRDS 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
-TABLE V: Comparison of CrackSCF Networks using SCFM, Concat and PAF. Results are highlighted with the best in red and the second-best in blue
+Original Conv 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
+DS Conv 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
+Low-Rank Conv 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
+LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
+TABLE V: Comparison of CrackSCF Networks using SCFM, signle Concat branch and signle PAF branch. Best results are bolded and second best results are underlined
 Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
-CrackSCF_Concat 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359 7.21G 4.71M 42
-CrackSCF_PAF 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427 8.24G 4.72M 39
-CrackSCF_SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
-TABLE VI: Comparison of results of CrackSCF networks trained with different loss ratios
-Dice : BCE ODS OIS P R F1 mIoU
-1 : 5 0.8129 0.8203 0.8217 0.8458 0.8336 0.8424
-1 : 4 0.8049 0.8145 0.8116 0.8365 0.8239 0.8368
-1 : 3 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
-1 : 2 0.8146 0.8211 0.8227 0.8435 0.8330 0.8438
-1 : 1 0.8136 0.8204 0.8231 0.8436 0.8332 0.8431
-Only BCE 0.8003 0.8056 0.8349 0.8004 0.8173 0.8332
-Only Dice 0.8107 0.8165 0.8302 0.8265 0.8283 0.8409
-IV-E Ablation Analysis
-To evaluate the advantages of our LRDS block compared to other convolution operations, we performed ablation experiments on the TUT dataset, as list in Table IV.
-On six metrics, our method slightly outperforms those using only depthwise separable or low-rank approximation convolutions.
-Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the low-rank approximation block, achieving the highest FPS.
+Signle Concat branch 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359 7.21G 4.71M 42
+Signle PAF branch 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427 8.24G 4.72M 39
+SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
+TABLE VI: Comparison of SCFM performance with other segmentation heads. Best results are bolded and second best results are underlined
+SegHead ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
+PID 0.7120 0.7198 0.7921 0.6742 0.7284 0.7777 36.69G 5.44M 21
+SegFormer 0.7446 0.7634 0.7533 0.7786 0.7658 0.7983 29.72G 6.20M 23
+UNet 0.7836 0.7951 0.7829 0.8248 0.8033 0.8222 40.59G 12.34M 20
+Hamburger 0.7447 0.7609 0.7075 0.8328 0.7670 0.7969 168.23G 7.45M 14
+SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
+TABLE VII: Comparison of the performance of using our lightweight deformable attention in LDE with other variants. Best results are bolded and second best results are underlined
+Attention ODS OIS P R F1 mIoU FLOPs Params FPS↑
+Common MHA 0.7919 0.8009 0.7999 0.8240 0.8118 0.8279 15.98G 5.80M 16
+MHA + LR_Linear 0.7924 0.8003 0.7988 0.8284 0.8134 0.8286 8.87G 4.77M 14
+Common MSDA 0.8058 0.8151 0.8052 0.8467 0.8257 0.8373 16.37G 5.82M 28
+MSDA + LR_Linear 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
+IV-D Complexity Analysis
+To evaluate our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS).
+FLOPs quantify computational load, reflecting demand during inference.
+Params measure storage requirements, reflecting trainable and stored parameters.
+FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
+Table III presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS.
+Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively.
+This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices.
+In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [8].
+This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications.
+Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
+IV-E Ablation Studies
+To evaluate the strengths of our proposed module over other modules, we performed ablation experiments on the TUT dataset.
+1) Ablation study of different convolution types: As list in Table IV, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution.
+Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS.
 Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance.
-These results indicate that our convolution block effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
-To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, we performed ablation experiments on the TUT dataset, as list in Table V.
+These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
+2) Ablation study in SCFM using different branch: To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, we performed ablation experiments on the TUT dataset, as list in Table V.
 Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics.
@@
 These results indicate that our module effectively balances performance and efficiency.
-Table VI shows the results of ablation experiments on different weight proportions of the two sub-losses in the loss function.
-As shown, a Dice loss to BCE loss ratio of 1:3 yields the highest values in five metrics: ODS, OIS, R, F1, and mIoU.
-Although the P metric is 0.80% lower than using only BCE loss, the F1 metric improves by 2.58%, indicating a better balance between P and R.
-Therefore, we use this ratio during training.
-In summary, a Dice loss to BCE loss ratio of 1:3 not only enhances pixel-level performance but also improves the global segmentation of crack areas, achieving superior segmentation performance.
-IV-F Limitations and Discussions
-Through these experiments, we found that SCFM effectively perceives and fuses local details and long-range pixel dependencies.
-Our network excels in various metrics on public datasets with simple scenes and performs exceptionally on the complex TUT dataset.
-The lightweight LRDS block minimizes parameters and computational load, facilitating deployment on resource-constrained devices like drones and smartphones.
-However, our network has several limitations, as follows:
-  1. 1.
-    Although our method generally achieves good crack segmentation with effective background noise suppression, as shown in the Fig.
-10, our model occasionally misdetects noise regions that resemble cracks in some images.
-It also fails to detect watermark occlusions with colors similar to the background.
-Therefore, it is essential to enhance modules such as MFE and LDE to improve the model’s capability to perceive and extract features from critical noise and occlusion areas, thereby further increasing its robustness and stability in complex scenarios.
-  2. 2.
-    Despite our method achieves the minimum parameter count and computational load, there is still room for improvement in inference speed.
-For example, our network achieves a processing speed of 36 frames per second, but the DTrCNet [9], which focuses more on inference speed, attains a faster speed of 47 frames per second.
-Therefore, it is necessary to further simplify convolution operations or the network’s complexity to achieve faster inference speeds while maintaining a good balance between segmentation performance and model lightweighting.
-Figure 10: Examples of CrackSCF misdetections and misses. The critical regions are marked with blue boxes.
+3) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [50], SegFormer [51], UNet [52], and Hamburger head [53], using the TUT dataset.
+As list in the Table VI, SCFM achieves the best results in segmentation performance metrics.
+Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively.
+This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels.
+Furthermore, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS.
+This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
+4) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA.
+As shown in the Table VII, our method achieves the best results in segmentation performance metrics.
+Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively.
+Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of our use of MSDA.
+Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease.
+For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively.
+This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
 V Conclusion
-This paper proposes a staircase cascaded fusion network (CrackSCF) for pixel-level crack detection.
-We developed SCFM to fuse the local patterns processed by MFE and the pixel dependencies processed by LDE.
-Additionally, we introduced the lightweight convolution block LRDS to replace all convolution operations in the network, effectively reducing the substantial parameter count and computational load required for training and inference.
+This paper proposes a staircase cascaded fusion crack segmentation network (CrackSCF) for pixel-level crack detection.
+In CrackSCF, we propose the lightweight convolutional block LRDS to construct the MFE, which dramatically reduces the amount of computation and parametric quantities required by the model and is able to efficiently capture the local details of the crack image, and we propose LDE to more efficiently perceive the pixel irregular long-range dependencies in the crack feature maps, in order to more comprehensively fuse the local patterns with the pixel’s long-range dependencies, we also propose SCFM to fully perceive them to generate higher quality segmentation maps.
 Furthermore, to comprehensively evaluate our model’s robustness and stability across various complex scenarios, we collected the TUT dataset, which includes eight types of crack scenes.
 Experimental results show that our method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation.
-The computational load and parameter count are reduced to 9.26G and 4.79M, respectively, with an inference speed of up to 36 frames per second.
-In future work, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.
-Additionally, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications.
+The computational load and parameter count are reduced to 9.26G and 4.79M, respectively.
+Despite our method achieves the low parameter count and computational load, there is still room for improvement in inference speed.
+In future work, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications.
+Additionally, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.

### v4

@@
-Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation
+CrackSCF: Lightweight Cascaded Fusion Network for Robust and Efficient Structural Crack Segmentation
 Abstract
-Detecting cracks with pixel-level precision for key structures is a significant challenge, existing methods struggle to integrate local textures and pixel dependencies of cracks.
-Furthermore, these methods possess numerous parameters and substantial computational requirements, complicating deployment on edge devices.
-In this paper, we propose the Staircase Cascaded Fusion Crack Segmentation Network (CrackSCF), which generates high-quality crack segmentation maps while reducing computational overhead.
-We design a lightweight convolutional block that substitutes all convolution operations, reducing the model’s computational demands while maintaining an effective capture of local details.
-Additionally, we introduce a lightweight long-range dependency extractor to better capture the long-range dependencies.
-Furthermore, we develop a staircase cascaded fusion module, which seamlessly integrates local patterns and long-range dependencies, resulting in high-quality segmentation maps.
+Accurately segmenting structural cracks at the pixel level remains a major hurdle, as existing methods fail to integrate local textures with pixel dependencies, often leading to fragmented and incomplete predictions.
+Moreover, their high parameter counts and substantial computational demands hinder practical deployment on resource-constrained edge devices.
+To address these challenges, we propose CrackSCF, a Lightweight Cascaded Fusion Crack Segmentation Network designed to achieve robust crack segmentation with exceptional computational efficiency.
+We design a lightweight convolutional block (LRDS) to replace all standard convolutions.
+This approach efficiently captures local patterns while operating with a minimal computational footprint.
+For a holistic perception of crack structures, a lightweight Long-range Dependency Extractor (LDE) captures global dependencies.
+These are then intelligently unified with local patterns by our Staircase Cascaded Fusion Module (SCFM), ensuring the final segmentation maps are both seamless in continuity and rich in fine-grained detail.
 To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets.
-The results show that our method outperforms existing methods, particularly in handling background noise and achieving detailed segmentation.
-The F1 and mIoU scores on the TUT dataset are 0.8382 and 0.8473, respectively, demonstrating state-of-the-art (SOTA) performance with low computational resources.
+The experimental results show that the CrackSCF method consistently outperforms the existing methods, and it demonstrates greater robustness in dealing with complex background noise.
+On the TUT dataset, CrackSCF achieved 0.8382 on F1 score and 0.8473 on mIoU, and it only required 4.79M parameters.
 The code and dataset is available at https://github.com/Karl1109/CrackSCF.
-Index Terms:
-Structural crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
-publicationid: pubid: $0$–$00$ 2025 IEEE
-I Introduction
-Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues.
+{IEEEkeywords}
+Crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
+1 Introduction
+[…]
+Cracks are common defects in industrial structures such as pavements and buildings, which can lead to serious safety issues.
 Thus, automated detection on edge control devices is crucial for ensuring production safety [1].
-Early automated detection methods used traditional image processing techniques for crack extraction [2, 3, 4].
+Early automated detection methods used traditional image processing techniques for crack extraction [2].
 While these methods are easy to implement, their performance in complex scenarios still needs improvement.
-Figure 1: Performance comparison of our proposed method with UCTNet [5], SFIAN [6], CT-crackseg [7], DTrCNet [8], Crackmer [9] and Simcrack [10] on the TUT dataset in terms of mIoUs, FLOPs and Params metrics.
-In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as MST-Net [11] and FcaHRNet [12] have shown efficient feature extraction capabilities in structural crack segmentation tasks.
+Figure 1: (a) shows the performance of our method compared to other methods. (b) shows the different scenarios included in TUT dataset.
+In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as Crack500 [3] and SC-OSDA [4] have shown efficient feature extraction capabilities in structural crack segmentation tasks.
 The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations.
@@
 Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
-The success of Vision Transformer (ViT) [13] and Swin Transformer [14] has demonstrated the advantages of Transformer [15], with their strong sequence processing capabilities, in handling visual tasks [16, 17].
-Unlike CNNs, transformers can capture dependencies between any pixels when processing images [18, 19], making them more advantageous in handling complex topologies.
-To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [20], MFAFNet [21], and Crackmer [8], which integrate local and long-range information through U-shaped connections or residual connections.
+The success of Vision Transformer (ViT) [5] has demonstrated the advantages of Transformer, with their strong sequence processing capabilities, in handling visual tasks [6].
+Unlike CNNs, transformers can capture dependencies between any pixels when processing images, making them more advantageous in handling complex topologies.
+To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as DTrCNet [7] and Crackmer [7], which integrate local and long-range information through U-shaped or residual connections.
 Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations, as follows:
   1. 1.
-    The different information fusion strategies significantly affect the segmentation performance of CNN-Transformer networks.
-For instance, UCTNet [5] and CT-crackseg [7] use channel concatenation to fuse local patterns and long-range dependences information.
+    The different fusion strategies significantly affect the segmentation performance of CNN-Transformer networks.
+For instance, UCTNet [8] and CTCS [9] use channel concatenation to fuse local patterns and long-range dependences information.
 However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise.
 Additionally, DTrCNet [7] and Crackmer [10] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance.
+  2. 2.
+    Most networks, in pursuit of performance, overlook the parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices.
 As shown in Fig.
-1(a), due to these shortcomings, these four methods fail to achieve the best performance.
-  2. 2.
-    Most networks, in pursuit of performance, overlook the significant parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices.
-As shown in Fig.
-1(b) and Fig.
-1(c), networks like SFIAN [6], Crackmer [9], and CT-crackseg[7] have very high computational FLOPs and parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
+1(a), networks like SFIAN [11], Crackmer [10], and CTCS[9] have high computational parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
 To address these issues, we propose a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise.
@@
 Existing public datasets like DeepCrack [12] and Crack500 [3] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance.
-The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines.
+As shown in Fig.
+1(b), the TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines.
 With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness.
@@
     We carried out comprehensive experiments on five public datasets and the TUT benchmark dataset, demonstrating that our proposed CrackSCF network outperforms existing SOTA methods.
-The organization of the subsequent sections of this paper is as follows: In Section II, existing crack detection models and lightweight networks are reviewed.
-Section III details the components of the proposed network.
-Section IV delves into the experimental details and results.
-Finally, Section V provides a summary of the entire paper and highlights future research directions.
-II Related Work
+The organization of the subsequent sections of this paper is as follows: In Section 2, existing crack detection models and lightweight networks are reviewed.
+Section 3 details the components of the proposed network.
+Section 4 delves into the experimental details and results.
+Finally, Section 5 provides a summary of the entire paper and highlights future research directions.
+2 Related Work
 Figure 2: General Architecture Diagram of CrackSCF Network.
@@
 Through four stages, resulting in the segmentation output.
-II-A Crack Segmentation Networks
+2.1 Crack Segmentation Networks
 Recently, a growing number of researchers have utilized CNNs to develop crack segmentation networks, achieving advanced performance.
-Al-Huda et al.
-[24] proposed the MDAUNet network, based on U-Net and dual attention modules, and designed a hybrid weighted segmentation loss function to address class imbalance issues.
+Zhu et al.
+[13] proposed the BSNBM network, which integrates bidirectional self-rectifying modules with Bayesian inference to enhance crack detection robustness and accuracy under uncertainty.
 Cheng et al.
 [11] introduced the Selective Feature Fusion and Irregular-Aware Network (SFIAN) to effectively model irregular crack objects, resulting in clearer backgrounds on various crack images.
-Huang et al.
-[25] developed a crack detection network for masonry structures that combines thermal imaging and visible light image fusion with semantic segmentation, laying the foundation for the development of cross-modal methods for structural crack monitoring.
-Although the aforementioned methods have demonstrated the advantages of CNNs in crack segmentation tasks, CNNs are essentially local operations and cannot effectively perceive a small number of pixels amidst a large background.
-Therefore, these methods often result in discontinuous crack pixels and are highly susceptible to background noise interference.
+Lyu et al.
+[14] proposed a CNN-based crack quantification framework that utilizes voxel-based reconstruction and Bayesian data fusion to achieve accurate crack width estimation.
 Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes.
 Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation.
-Chen et al.
-[26] proposed TransUNet, a novel network that effectively integrates long-range information and local features through a U-Net-like structure.
 Qi et al.
@@
 [16] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
-The aforementioned methods, while demonstrating the advantages of CNN-Transformer networks in processing pixel sequences and suppressing background noise, have limitations due to using only one method or branch for feature fusion.
-This reduces their ability to effectively perceive long-range dependencies and local patterns, causing some complex regions’ local or long-range information to be overshadowed.
-Additionally, these methods excessively focus on performance, overlooking the substantial computational resources required for feature extraction and fusion, making them difficult to deploy on edge computing devices.
-II-B Lightweight Networks
+Although CNN and CNN-Transformer methods have shown promise in crack segmentation, CNNs struggle to capture global context due to their local receptive fields, leading to discontinuous predictions and vulnerability to background noise.
+CNN-Transformer hybrids improve global perception but often adopt single-branch fusion, limiting their ability to integrate local and long-range cues.
+Moreover, both lack selective interaction and noise suppression, and their high computational cost hinders deployment on edge devices.
+2.2 Lightweight Networks
 Deploying deep learning models on edge devices like smartphones and drones is trending.
 However, these devices have limited computing resources, making it challenging to use models with high parameter counts.
-Thus, compact model design for lightweight modifications in size and speed is crucial [29, 30].
+Thus, compact model design for lightweight modifications in size and speed is crucial.
 For example, Howard et al.
@@
 [21] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection.
-Although these lightweight models reduce computational resources, their emphasis on minimizing parameters results in limited model expressiveness, hindering their ability to effectively capture intricate patterns and features in the input data.
-This leads to unsatisfactory segmentation results.
-This imbalance between performance and lightweight design makes them unreliable in practical applications.
-III Methodology
-III-A Network Architecture
+Although lightweight models reduce computational cost, their aggressive parameter reduction limits model expressiveness, impairing the capture of intricate patterns and leading to suboptimal segmentation.
+This trade-off between performance and efficiency undermines their reliability in real-world applications.
+3 Methodology
+3.1 Network Architecture
 As shown in Fig.
@@
 In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
-Specifically, for a given dataset, the crack images are first input into the MFE module, which is divided into four stages.
+Specifically, for a given dataset, the images are first input into the MFE, which is divided into four stages.
 In each stage, the extracted feature maps undergo a Feature Enhancement stage.
-Then, the LDE module processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies.
-Subsequently, the SCFM module receives these sequences along with the four layers of feature maps extracted by the MFE.
+Then, the LDE processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies.
+Subsequently, the SCFM receives these sequences along with the four layers of feature maps extracted by the MFE.
 After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
@@
 Replacing all convolutions with LRDS blocks significantly reduced FLOPs and Params, demonstrating the effectiveness of LRDS in model lightweighting.
-Figure 3: Comparison of FLOPs and Params for each module using the original convolution and our proposed LRDS convolution block.
-III-B Lightweight Multi-scale Feature Extractor
+Figure 3: Comparison of each module when using our LRDS or original convolution separately. (a) shows the comparison of FLOPs, (b) shows the comparison of parameters.
+3.2 Lightweight Multi-scale Feature Extractor
 As shown in Fig.
-4, our designed LRDS block adopts a bottleneck-like structure.
+2, our designed LRDS block adopts a bottleneck-like structure.
 It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer.
-The input feature map first passes through the Reduce layer, reducing the input channels to $C_{m}$.
+The input feature map first passes through the Reduce layer, reducing the input channels to $f_{mid}$.
 It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged.
-Finally, the Expand layer increases the number of channels to $C_{d}$, producing the final output feature map.
+Finally, the Expand layer increases the number of channels to $f_{out}$, producing the final output feature map.
 Low-rank approximation decomposes a high-dimensional matrix into low-dimensional matrices, reducing computation complexity.
@@
 Given that $O(ee_{0})\ll O(e_{0}k^{2}c)$, the complexity of using Equation (3) is $\frac{e_{0}}{e}$ times that of using Equation (1).
-Low-rank approximation theory greatly reduces the convolutional layer parameters and computational load.
 In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space.
-We constructed MFE by replacing all convolution operations in ResNet50 with LRDS convolution blocks.
-Figure 4: Illustration of LRDS convolution block.
-After the number of channels of the input features is reduced, their local and spatial features are efficiently extracted by depth-wise convolution and point-wise convolution, and finally the number of channels is restored.
-III-C Lightweight Long-range Dependency Extractor
+We constructed MFE by replacing all convolution operations with LRDS convolution blocks.
+3.3 Lightweight Long-range Dependency Extractor
 Figure 4: Analysis of Computational Resources Required for Linear Operations.
@@
 Specifically, given an input $x\in\mathbb{R}^{C\times H\times W}$, let $q$ denote a query element with query feature $z_{q}$ and reference point coordinates $(p_{qx},p_{qy})$, the deformable attention is calculated as:
-$\centering\begin{split}&\text{MSDeformAttn}(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg{[}\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg{]}\end{split}\@add@centering$ (4)
+$\centering\begin{split}&MSDeformAttn(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg[\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg]\end{split}\@add@centering$ (4)
 where $h$ represents the attention head, $c$ represents the input feature layer, and $t$ represents the sampling key point.
@@
 As shown in Fig.
-5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module.
+4(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module.
 In order to reduce the computational resources required by the LDE, we propose a linear transformation method (LR Linear) based on low-rank approximation.
@@
 As shown in Fig.
-5(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
+4(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
 The structure of LDE is shown in Fig.
@@
 The encoder then combines LR Linear with lightweight deformable attention mechanisms to capture long-range dependencies:
-$Z=\text{MSDeformAttn}(S,P,S)$ (8)
+$Z=MSDeformAttn(S,P,S)$ (8)
 where $P\in\mathbb{R}^{B\times L\times 2}$ represent the normalized reference coordinates.
 Subsequently, $Z$ is processed through a feedforward network (FFN) that combines LR Linear to refine the features:
-$M=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Z)))$ (9)
+M=LR_Linear(GELU(LR_Linear(Z)))M=LR\_Linear(GELU(LR\_Linear(Z))) (9)
 where the output memory matrix $M\in\mathbb{R}^{B\times L\times d}$ is used as the input to the decoder, encoding multi-scale local dependencies.
@@
 First, the reference points $H\in\mathbb{R}^{N\times 2}$ are mapped to the encoded memory $M$ to aggregate features:
-$Y_{0}=\text{MSDeformAttn}(Q,H,M)$ (10)
+$Y_{0}=MSDeformAttn(Q,H,M)$ (10)
 Then, it passes through the FFN to obtain the crack pixel sequence enriched with irregular long-range dependencies:
-$Y=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Y_{0})))$ (11)
+$Y=LR\_Linear(GELU(LR\_Linear(Y_{0})))$ (11)
 where Y is the output sequence of the LDE enriched with irregular long-range dependencies between pixels.
-III-D Staircase Cascaded Fusion Module
-Figure 6: Illustration of the PAF. It effectively reduces required computing resources by embedding LRDS.
+3.4 Staircase Cascaded Fusion Module
 The design of our SCFM structure is shown in Fig.
@@
 The structure of PAF is shown in Fig.
-6.
+2.
 We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively.
 The output can be expressed as:
-$\sigma=\text{Sigmoid}(f_{c}(\vec{v_{c}})\cdot f_{d}(\vec{v_{d}}))$ (12)
-where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. Thus, the output of PAF can be represented as:
-$\text{Output}=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (13)
+$Output=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (12)
 where $\sigma$ indicates the probability that the two pixels belong to the same category.
@@
 The processes of SCFM can be given by the following equation:
-$\begin{split}&F_{i}^{\text{O}}=C_{\text{i}}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\right)\odot P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\end{split}$ (14)
+$\begin{split}&F_{i}^{O}=C_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\right)\odot P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\end{split}$ (13)
 Where $F_{i}^{\text{O}}$ represents the output features of the $i$-th SCFM layer ($i$ ranges from 0 to 3), $F_{i}^{\text{MFE}}$ represents the $i$-th layer MFE features, $C_{\text{i}}$ is an operation chain: Concat followed by LRDS, ReLU, and GN, $P_{i}$ is a PAF operation, $\mu(\cdot)$ is the sigmoid function generating spatial attention weights, and $\odot$ represents the Hadamard product.
@@
 Furthermore, the fine-grained features extracted with PAF can be perceived more finely through the dynamic weighting mechanism to achieve a clearer segmentation map.
-III-E Loss Function
+3.5 Loss Function
 The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance.
-Previously, Binary Cross-Entropy (BCE) [36] was often used, but it may not effectively distinguish the minority crack pixels in the image.
-Combining BCE with Dice loss [37] addresses this issue.
-BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data.
+Previously, Binary Cross-Entropy (BCE) was often used, but it may not effectively distinguish the minority crack pixels in the image.
+Combining BCE with Dice loss addresses this issue.
+BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data [22] .
 Our loss $L$ can be calculated using Equation (14), which combines the above two loss functions for joint optimization.
@@
 where $\mathit{L_{B}}$ represents the BCE loss function and $\mathit{L_{D}}$ represents the Dice loss function.
-IV Experiments
+4 Experiments
 In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics.
 We then analyze results and model complexity, perform ablation experiments.
-IV-A Dataset Description
-TABLE I: Comparison of Different Crack Datasets
+4.1 Dataset Description
+Table 1: Comparison of Different Crack Datasets
 No. Datasets Year Num(train/val/test) Resolution Percentage of cracked pixels Scenarios
@@
 5 CrackMap [25] 2023 120(72/12/36) 256 width and 256 height 4.10% bitumen
-6 AEL [23] 2019 58 768 width and 512 height 0.67% bitumen
-7 CrackTree206 [41] 2012 206 800 width and 600 height 0.32% bitumen
-8 LSCD [42] 2024 143 400 width and 400 height 3.18% bitumen, cement
-9 CrackSC [43] 2023 197 320 width and 480 height 2.02% bitumen, cement
 6 TUT (Ours) 2024 1408(987/139/282) 640 width and 640 height 3.16% bitumen, cement, bricks, runways, metal, blades, pipelines, tiles
-Figure 7: Illustration of the images of each scenarios in the TUT dataset.
 The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only.
@@
 To tackle this problem, we created the TUT dataset to enable better segmentation across different scenarios.
-Detailed information and comparisons with other datasets are list in Table I.
-The TUT dataset comprises 1270 crack images captured with our mobile phones, along with a small number of 138 images sourced from the Internet.
+Detailed information and comparisons with other datasets are list in Table 1.
+The TUT dataset comprises 1270 crack images, along with a small number of 138 images sourced from the Internet.
 TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances.
 To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images.
-Unlike other datasets with clear backgrounds, the TUT dataset images have complex, noisy backgrounds and cracks of intricate shapes.
+Unlike other datasets with clear backgrounds, the TUT dataset images have complex backgrounds and cracks of intricate shapes.
 This evaluates the model’s capability to segment cracks in challenging conditions.
-As list in Table I, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios.
+Figure 5: Illustration of the various scenarios of the TUT dataset with the original images and Ground Truth.
+As list in Table 1, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios.
 Fig.
@@
 As seen in Fig.
-7, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties.
+5, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties.
 Particularly in plastic runways, both coarse and fine cracks are present, testing the model’s generalization to different types of cracks.
@@
 This can easily result in false detections, thereby testing the model’s ability to accurately extract crack areas from complex environments.
-IV-B Implementation Details
+4.2 Implementation Details
 1) Experimental Settings: The CrackSCF network is built on the PyTorch v2.1.1 deep learning framework and trained on an Ubuntu 20.04 server, featuring two Intel(R) Xeon(R) Platinum 8336C CPUs and eight Nvidia GeForce 4090 GPUs, each with 24GB VRAM.
@@
 We compared the CrackSCF network with other SOTA methods on six datasets: CFD [23], GAPS509 [24], DeepCrack [12], Crack500 [3], CrackMap [25], and ours TUT dataset.
-2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation methods and the latest SOTA methods on different datasets.
-These methods include HED [44], UNet++ [45], Deeplabv3+ [46], AttuNet [47], RCF [48], RIND [49], UCTNet [5], SFIAN [6], CT-crackseg [7], DTrCNet [8], Crackmer [9] and Simcrack [10].
-TABLE II: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
+Table 2: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
 Methods CFD GAPS509 Crack500
 ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU
-HED 0.6276 0.6494 0.5467 0.7374 0.6279 0.7240 0.4014 0.4509 0.4300 0.4624 0.4456 0.6282 0.6191 0.6537 0.6150 0.7281 0.6668 0.7140
-UNet++ 0.6586 0.6836 0.5995 0.7608 0.6706 0.7432 0.6255 0.6484 0.6160 0.6684 0.6411 0.7300 0.6965 0.7274 0.6989 0.7805 0.7374 0.7605
-DeeplabV3+ 0.6910 0.7084 0.6264 0.7730 0.6920 0.7709 0.6160 0.6425 0.6050 0.6669 0.6345 0.7378 0.6943 0.7211 0.6970 0.7569 0.7257 0.7746
-AttuNet 0.6710 0.6870 0.6240 0.7220 0.6694 0.7576 0.5955 0.6174 0.5924 0.6519 0.6208 0.7247 0.6586 0.6890 0.6490 0.7608 0.7005 0.7543
-RCF 0.6434 0.6624 0.5792 0.7348 0.6478 0.7343 0.4532 0.4817 0.4420 0.5428 0.4872 0.6472 0.6054 0.6361 0.6037 0.7136 0.6540 0.7068
 RIND 0.6858 0.7021 0.6456 0.7534 0.6954 0.7592 0.5505 0.6023 0.4959 0.6494 0.5623 0.6909 0.6469 0.6483 0.6998 0.7245 0.7119 0.7381
@@
 SFIAN 0.7040 0.7159 0.6657 0.7512 0.7058 0.7684 0.6087 0.6320 0.6061 0.6451 0.6250 0.7198 0.6977 0.7348 0.6983 0.7742 0.7343 0.7604
-CT-crackseg 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
+CTCS 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
 DTrCNet 0.7079 0.7123 0.6774 0.7486 0.7112 0.7710 0.6211 0.6276 0.6277 0.6592 0.6431 0.7272 0.7012 0.7241 0.6527 0.8280 0.7357 0.7627
@@
 CrackMap DeepCrack TUT (Ours)
-HED 0.6022 0.6240 0.5281 0.7063 0.6043 0.7026 0.7552 0.8237 0.7231 0.7678 0.7448 0.8007 0.6960 0.7254 0.7080 0.7431 0.7251 0.7683
-UNet++ 0.7030 0.7222 0.6773 0.7209 0.6984 0.7631 0.8419 0.8893 0.8474 0.8533 0.8503 0.8621 0.7702 0.7948 0.7725 0.7974 0.7848 0.8160
-DeeplabV3+ 0.6617 0.6793 0.5853 0.7532 0.6587 0.7460 0.7727 0.8354 0.7363 0.8323 0.7813 0.8565 0.7504 0.7777 0.7155 0.8231 0.7655 0.8177
-AttuNet 0.6618 0.6773 0.6233 0.6970 0.6581 0.7454 0.8153 0.8379 0.8066 0.8069 0.8068 0.8578 0.7539 0.7795 0.7558 0.7945 0.7747 0.8204
-RCF 0.5552 0.5683 0.4590 0.7164 0.5595 0.6758 0.7666 0.8376 0.7423 0.7611 0.7516 0.8110 0.6821 0.7138 0.7106 0.6877 0.6990 0.7598
 RIND 0.6745 0.6943 0.6023 0.7586 0.6699 0.7425 0.8087 0.8267 0.7896 0.8920 0.8377 0.8391 0.7531 0.7891 0.7872 0.7665 0.7767 0.8051
@@
 SFIAN 0.7200 0.7465 0.6715 0.7668 0.7160 0.7748 0.8616 0.8928 0.8549 0.8692 0.8620 0.8776 0.7290 0.7513 0.7715 0.7367 0.7537 0.7896
-CT-crackseg 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
+CTCS 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
 DTrCNet 0.7328 0.7413 0.6912 0.7681 0.7276 0.7812 0.8473 0.8512 0.8905 0.8251 0.8566 0.8661 0.7987 0.8073 0.7972 0.8441 0.8202 0.8331
@@
 Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
-3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), Mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
-Mean Intersection over Union (mIoU) measures the average ratio of the intersection and union of the predicted results and the true results.
-The calculation formula for mIoU is:
-$MIoU=\frac{1}{i+1}\sum_{a=0}^{i}\frac{p_{aa}}{\sum_{b=0}^{i}p_{ab}+\sum_{b=0}^{i}p_{ba}-p_{aa}}$ (16)
-where $i$ is the number of classes, $a$ denotes the true value, $b$ denotes the predicted value, and $p_{ab}$ represents the number of pixels predicted as $a$ for $b$.
-Given that our crack segmentation task involves binary classification, we set $i=1$.
-Optimal Dataset Scale (ODS) indicates the optimal segmentation performance metric obtained when a fixed threshold $t$ is selected over the entire dataset. It is defined as:
-$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (17)
-Optimal Image Scale (OIS) indicates the optimal segmentation performance metric obtained when an optimal threshold $t$ is selected for each image individually. It is defined as:
-$OIS=\frac{1}{N}\sum_{a=1}^{N}\max_{t}\frac{2\cdot P_{t,a}\cdot R_{t,a}}{P_{t,a}+R_{t,a}}$ (18)
-where $N$ represents the total number of images.
-IV-C Comparison with the SOTA methods
-Our proposed method, CrackSCF, was compared with 12 other methods across five public datasets and our own TUT dataset.
-Table II displays the comparison results for the five public datasets, while Fig.
-8 illustrates the visual results of all methods on these datasets.
-Figure 8: Visualization results of CrackSCF and other methods on public datasets.
-Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
-1) Results on the CFD dataset: The crack areas in the CFD dataset are very slender, testing the model’s sensitivity to fine-grained crack pixels.
-As list in Table II, our proposed CrackSCF attained the best results across all metrics for the CFD dataset.
-Specifically, on ODS, OIS, P, R, F1, and mIoU, it outperformed the second-best methods by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively.
-Notably, SimCrack [10] achieved the second-best performance on all metrics except for P and R, only behind our method.
-However, as seen in Fig.
-8, its actual segmentation performance is not satisfactory, while our method achieves high-quality segmentation results on various fine-grained crack pixels.
-Other methods, such as UCTNet [5] and DTrCNet [8], tend to misdetect background noise or miss key detail areas of crack pixels.
-2) Results on the GAPS509 dataset: Unlike the CFD dataset, the GAPS509 dataset contains not only finer cracks but also some irrelevant background interference, testing the model’s ability to extract crack areas under high interference conditions.
-As list in Table II, our method achieved the highest performance on all metrics except R.
-Although SimCrack [10] achieved 0.6997 on the R metric, leading our method by about 2.72%, its P metric was 0.6171, far lower than our method.
-The highest F1 score obtained by our method indicates a good balance between P and R.
-Specifically, our method outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.31%, 1.25%, 2.90%, 1.11%, and 0.45%, respectively.
-Additionally, as seen in Fig.
-8, our method almost perfectly avoids irrelevant background noise, while most other methods falsely detect irrelevant backgrounds and perform worse in segmenting crack pixel areas compared to our method.
-3) Results on the DeepCrack dataset: The cracks in the DeepCrack dataset are often relatively wide, testing the model’s ability to extract a large range of crack pixels.
-As list in Table II, our method achieved the highest results on all metrics for the DeepCrack dataset.
-Specifically, our method outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85%, respectively.
-Notably, CT-crackseg also achieved good results on this dataset, but as seen in Fig.
-8, its segmentation performance on detailed cracks is not as good as our method, indicating that our method can effectively handle crack details.
-4) Results on the Crack500 dataset: Similar to the DeepCrack, the Crack500 dataset also has relatively wide crack areas, but its structure is complex, posing higher requirements on model performance.
-As list in Table II, our method achieved the best results on five metrics for the Crack500 dataset, except for R.
-It outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%, respectively.
-Although DTrCNet [8] leads on the R metric, our method surpasses it on the P metric by 10.51%.
-The highest F1 score achieved by our method indicates a good balance between P and R.
-Additionally, as seen in Fig.
-8, our method does not produce false detections in irrelevant background areas, whereas other methods either produce false detections or lack detailed segmentation in crack areas.
-5) Results on the CrackMap dataset: Similar to the CFD dataset, the cracks in the CrackMap dataset are also relatively slender.
-As list in Table II, our method leads all other methods on all metrics, outperforming the second-best methods on ODS, OIS, P, R, F1, and mIoU by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively.
-Notably, SimCrack [10] achieved the second-best results on this dataset, but as seen in Fig.
-8, SimCrack’s segmentation performance on the crack tips is not satisfactory, whereas our method effectively identifies these pixels.
-Figure 9: Visualization results of CrackSCF and other methods on TUT dataset.
-Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
-TABLE III: Comparison of the complexity of CrackSCF with other methods. Best results are bolded and second best results are underlined
-Methods FLOPs ↓ Params ↓ FPS ↑
-HED 80.35G 14.72M 33
-RCF 115.54G 15.52M 30
-RIND 695.77G 59.39M 11
-UCTNet 171.77G 66.22M 12
-UNet++ 139.61G 9.16M 22
-DeeplabV3+ 88.53G 59.34M 29
-AttuNet 541.34G 57.16M 24
-SFIAN 52.36G 13.63M 35
-CT-crackseg 39.47G 22.88M 28
-DTrCNet 123.20G 63.45M 47
-SimCrack 70.66G 29.58M 29
-Crackmer 14.94G 5.90M 31
-Ours 9.26G 4.79M 36
-6) Results on the TUT dataset: Fig.
+2) Comparison Methods: To evaluate the model performance, we compared the CrackSCF with seven classical and latest SOTA methods on different datasets.
+These methods include RIND [26], UCTNet [8], SFIAN [11], CTCS [9], DTrCNet [7], Crackmer [10] and Simcrack [27].
+3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics [11]: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
+4.3 Comparison with the SOTA methods
+1) Results on the our TUT dataset: Fig.
 6 shows the visual results of all methods on the TUT dataset.
 We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes.
-As list in Table II, our method achieved excellent results on the TUT dataset, leading in all metrics.
+As list in Table 2, our method achieved excellent results on the TUT dataset, leading in all metrics.
 Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively.
 SimCrack [27] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig.
-9, its segmentation performance on some key details has room for improvement compared to our method.
+6, its segmentation performance on some key details has room for improvement compared to our method.
 Furthermore, most methods fail to suppress irrelevant background noise.
@@
 Our method demonstrates excellent robustness and versatility, resulting in high-quality crack segmentation maps and adapting well to different scenarios.
-TABLE IV: Comparison of CrackSCF using LRDS blocks with networks using other convolutions. Best results are bolded and second best results are underlined
+2) Results on other public datasets: The CFD, CrackMap and GAPS509 datasets are characterized by slender and fine-grained crack structures, testing the model’s sensitivity to subtle and narrow crack pixels.
+As listed in Table 2, our proposed CrackSCF achieved the best performance on the CFD and CrackMap datasets.
+On both datasets, CrackSCF outperforms Simcrack by an average of 1.97%, 9.89%, 2.85%, 1.83%, and 1.00% on ODS, OIS, P, F1, and mIoU respectively.
+Although SimCrack [27] ranked second on most metrics, its actual segmentation performance on fine crack pixels and crack tips was unsatisfactory, as shown in Fig.
+6.
+Other methods such as UCTNet [8] and DTrCNet [7] often failed to distinguish crack regions from background noise or missed crucial fine-grained crack details.
+On the more challenging GAPS509 dataset, our method also achieved the highest performance on all metrics except recall.
+SimCrack obtained a higher recall value of 0.6997, about 2.72% higher than our method, but its precision was significantly lower at 0.6171, leading to a lower F1 score overall.
+Our method maintained a superior balance between precision and recall, with improvements of 1.31%, 1.25%, 2.90%, 1.11%, and 0.45% on ODS, OIS, P, F1, and mIoU, respectively.
+Visual results in Fig.
+6 further confirm that our method effectively suppresses false detections in irrelevant background areas while preserving accurate segmentation of slender cracks.
+Figure 6: Visualization results of CrackSCF and other methods on TUT and other public dataset.
+Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
+Both DeepCrack and Crack500 datasets are characterized by relatively wide crack regions, testing the model’s capability to capture extensive crack areas.
+As listed in Table 2, our method achieved the best results across all metrics on the DeepCrack dataset, outperforming the second-best methods by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85% in terms of ODS, OIS, P, R, F1, and mIoU.
+Although CTCS [9] also performed well, it fell short in preserving crack detail segmentation, as illustrated in Fig.
+6.
+This demonstrates the advantage of our method in accurately segmenting both wide crack regions and their finer structures.
+On the Crack500 dataset, our method achieved leading performance in five of the six evaluation metrics: ODS, OIS, P, F1, and mIoU, surpassing the second-best methods by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%.
+While DTrCNet [7] achieved a higher recall, our method outperformed it in precision by a notable 10.51%, leading to the highest F1 score and indicating a superior balance between detecting true crack pixels and suppressing false positives.
+Furthermore, as shown in Fig.
+6, our method produces fewer erroneous detections in irrelevant background areas and achieves more accurate and refined segmentation of complex crack structures compared to competing approaches.
+Table 3: Comparison of the complexity of CrackSCF with other methods
+Methods FLOPs ↓ Params ↓ FPS ↑
+RIND 695.77G 59.39M 11
+UCTNet 171.77G 66.22M 12
+SFIAN 52.36G 13.63M 35
+CTCS 39.47G 22.88M 28
+DTrCNet 123.20G 63.45M 47
+Crackmer 14.94G 5.90M 31
+SimCrack 70.66G 29.58M 29
+Ours 9.26G 4.79M 36
+Table 4: Comparison of CrackSCF using LRDS blocks with networks using other convolutions
 Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
@@
 LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
-TABLE V: Comparison of CrackSCF Networks using SCFM, signle Concat branch and signle PAF branch. Best results are bolded and second best results are underlined
-Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
-Signle Concat branch 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359 7.21G 4.71M 42
-Signle PAF branch 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427 8.24G 4.72M 39
-SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
-TABLE VI: Comparison of SCFM performance with other segmentation heads. Best results are bolded and second best results are underlined
+Table 5: Comparison of SCFM performance with other segmentation heads
 SegHead ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
@@
 SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
-TABLE VII: Comparison of the performance of using our lightweight deformable attention in LDE with other variants. Best results are bolded and second best results are underlined
+Table 6: Comparison of the performance of using our lightweight deformable attention in LDE with other variants
 Attention ODS OIS P R F1 mIoU FLOPs Params FPS↑
@@
 MSDA + LR_Linear 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
-IV-D Complexity Analysis
+4.4 Complexity Analysis
 To evaluate our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS).
@@
 FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
-Table III presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS.
-Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively.
+Table 3 presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS.
+Our approach achieves notably low values in FLOPs at 9.26G and Params at 4.79M, surpassing the second-best by 38.01% and 18.81%, respectively.
 This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices.
@@
 Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
-IV-E Ablation Studies
-To evaluate the strengths of our proposed module over other modules, we performed ablation experiments on the TUT dataset.
-1) Ablation study of different convolution types: As list in Table IV, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution.
+4.5 Ablation Studies
+We conducted ablation studies on the TUT dataset to assess the effectiveness of our proposed module.
+1) Ablation study of different convolution types: As list in Table 4, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution.
 Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS.
@@
 These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
-2) Ablation study in SCFM using different branch: To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, we performed ablation experiments on the TUT dataset, as list in Table V.
-Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics.
-Using only the PAF branch results in a 1.06% drop in the F1 score.
-While SCFM slightly increases parameter count, computational load, and inference speed, it significantly improves segmentation performance.
-These results indicate that our module effectively balances performance and efficiency.
-3) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [50], SegFormer [51], UNet [52], and Hamburger head [53], using the TUT dataset.
-As list in the Table VI, SCFM achieves the best results in segmentation performance metrics.
+2) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [28], SegFormer [6], UNet [29], and Hamburger head [30], using the TUT dataset.
+As list in the Table 5, SCFM achieves the best results in segmentation performance metrics.
 Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively.
@@
 This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
-4) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA.
-As shown in the Table VII, our method achieves the best results in segmentation performance metrics.
+3) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA.
+As shown in the Table 6, our method achieves the best results in segmentation performance metrics.
 Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively.
@@
 This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
-V Conclusion
+5 Conclusion
 This paper proposes a staircase cascaded fusion crack segmentation network (CrackSCF) for pixel-level crack detection.

### v5

@@
-CrackSCF: Lightweight Cascaded Fusion Network for Robust and Efficient Structural Crack Segmentation
+Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation
 Abstract
@@
 These are then intelligently unified with local patterns by our Staircase Cascaded Fusion Module (SCFM), ensuring the final segmentation maps are both seamless in continuity and rich in fine-grained detail.
-To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets.
+To comprehensively evaluate our method, this paper created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets.
 The experimental results show that the CrackSCF method consistently outperforms the existing methods, and it demonstrates greater robustness in dealing with complex background noise.
 On the TUT dataset, CrackSCF achieved 0.8382 on F1 score and 0.8473 on mIoU, and it only required 4.79M parameters.
-The code and dataset is available at https://github.com/Karl1109/CrackSCF.
-{IEEEkeywords}
-Crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
+keywords:
+Civil Infrastructure , Structural Crack , Lightweight Networks , Cascaded Fusion , Local Patterns , Long-Range Dependencies
+journal: Engineering Structures
+[…]
+[…]
+[…]
+[…]
+{graphicalabstract}[Uncaptioned image]
+{highlights}
+Research highlight 1
+This paper proposed a novel CrackSCF network for structural crack segmentation.
+By introducing a lightweight convolution block, a lightweight long-range dependency extractor, and a staircase cascaded fusion module, this network effectively distinguishes between background and crack pixels with low computational resources, resulting in high-quality segmentation maps.
+Research highlight 2
+This paper collected the TUT dataset, which includes structural cracks in eight different scenarios, providing a more comprehensive evaluation of the network’s performance.
+Research highlight 3
+Comprehensive experiments conducted on five public datasets and the TUT benchmark dataset demonstrate that the proposed CrackSCF network outperforms existing SOTA methods.
 1 Introduction
-[…]
-Cracks are common defects in industrial structures such as pavements and buildings, which can lead to serious safety issues.
-Thus, automated detection on edge control devices is crucial for ensuring production safety [1].
-Early automated detection methods used traditional image processing techniques for crack extraction [2].
+Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues.
+Thus, automated detection on edge control devices is crucial for ensuring production safety [1, 3, 26, 18, 21].
+Early automated detection methods used traditional image processing techniques for crack extraction [43, 39, 23, 31].
 While these methods are easy to implement, their performance in complex scenarios still needs improvement.
-Figure 1: (a) shows the performance of our method compared to other methods. (b) shows the different scenarios included in TUT dataset.
-In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as Crack500 [3] and SC-OSDA [4] have shown efficient feature extraction capabilities in structural crack segmentation tasks.
+In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as MST-Net [25] and FcaHRNet [5] have shown efficient feature extraction capabilities in structural crack segmentation tasks.
 The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations.
@@
 Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
-The success of Vision Transformer (ViT) [5] has demonstrated the advantages of Transformer, with their strong sequence processing capabilities, in handling visual tasks [6].
-Unlike CNNs, transformers can capture dependencies between any pixels when processing images, making them more advantageous in handling complex topologies.
-To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as DTrCNet [7] and Crackmer [7], which integrate local and long-range information through U-shaped or residual connections.
-Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations, as follows:
-  1. 1.
-    The different fusion strategies significantly affect the segmentation performance of CNN-Transformer networks.
-For instance, UCTNet [8] and CTCS [9] use channel concatenation to fuse local patterns and long-range dependences information.
+The success of Vision Transformer (ViT) [8] and Swin Transformer [20] has demonstrated the advantages of Transformer [32], with their strong sequence processing capabilities, in handling visual tasks [42, 45].
+Unlike CNNs, transformers can capture dependencies between any pixels when processing images [10, 29], making them more advantageous in handling complex topologies.
+To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [6], MFAFNet [7], and Crackmer [35], which integrate local and long-range information through U-shaped connections or residual connections.
+Figure 1: Performance comparison of our proposed method with UCTNet [33], SFIAN [4], CT-crackseg [30], DTrCNet [35], Crackmer [34] and Simcrack [13] on the TUT dataset in terms of mIoUs, FLOPs and Params metrics.
+Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations.
+The different information fusion strategies significantly affect the segmentation performance of CNN-Transformer networks.
+For instance, UCTNet [33] and CT-crackseg [30] use channel concatenation to fuse local patterns and long-range dependences information.
 However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise.
 Additionally, DTrCNet [35] and Crackmer [34] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance.
-  2. 2.
-    Most networks, in pursuit of performance, overlook the parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices.
-As shown in Fig.
-1(a), networks like SFIAN [11], Crackmer [10], and CTCS[9] have high computational parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
-To address these issues, we propose a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise.
-Specifically, to reduce the number of parameters and computational burden, we introduce a lightweight convolution block called LRDS.
-This module utilizes the principle of low-rank approximation to effectively reduce the computational resources required by convolutional layers, and captures spatial structures efficiently in the low-rank space through depthwise convolution, maintaining good feature perception while requiring very low computational resources.
-We also propose a lightweight Long-range Dependency Extractor (LDE), which uses a lightweight deformable attention mechanism to effectively capture irregular long-range dependencies between crack pixels with very low computational demands.
-Additionally, to better leverage the complementarity and correlation of feature maps while ensuring that critical detail information is not missed, we introduce the Staircase Cascaded Fusion Module (SCFM).
-This module generates feature maps layer by layer and effectively fuses them using a pixel attention mechanism that can finely capture local and long-range information and a channel concatenation mechanism, comprehensively combining local texture information and pixel dependencies.
-To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images.
+As shown in Figure 1(a), due to these shortcomings, these four methods fail to achieve the best performance.
+Most networks, in pursuit of performance, overlook the significant parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices.
+As shown in Figure 1(b) and Figure 1(c), networks like SFIAN [4], Crackmer [34], and CT-crackseg[30] have very high computational FLOPs and parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
+To address these issues, this paper proposes a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise.
+Specifically, to reduce the number of parameters and computational burden, a lightweight convolution block called LRDS is introduced.
+This module utilizes the principle of low-rank approximation to effectively reduce the computational resources required by convolutional layers and captures spatial structures efficiently in the low-rank space through depthwise convolution, maintaining strong feature perception.
+The proposed architecture also includes a lightweight Long-range Dependency Extractor (LDE), which uses a lightweight deformable attention mechanism to effectively capture irregular long-range dependencies between crack pixels with minimal computational demands.
+Additionally, to better leverage the complementarity and correlation of feature maps while ensuring that critical detail information is not missed, the Staircase Cascaded Fusion Module (SCFM) is incorporated.
+This module generates feature maps layer by layer and effectively fuses them using a pixel attention mechanism that finely captures local and long-range information along with a channel concatenation mechanism, comprehensively combining local texture information and pixel dependencies.
+To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, this paper also collected a dataset named TUT, which contains 1408 RGB images.
 Existing public datasets like DeepCrack [17] and Crack500 [40] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance.
-As shown in Fig.
-1(b), the TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines.
+The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines.
 With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness.
 Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
-In conclusion, our main contributions are as follows:
-  1. 1.
-    We proposed a novel CrackSCF network for structural crack segmentation.
-By introducing a lightweight convolution block, a lightweight long-range dependency extractor, and a staircase cascaded fusion module, this network effectively distinguishes between background and crack pixels with low computational resources, resulting in high-quality segmentation maps.
-  2. 2.
-    We collected the TUT dataset, which includes structural cracks in eight different scenarios, providing a more comprehensive evaluation of the network’s performance.
-  3. 3.
-    We carried out comprehensive experiments on five public datasets and the TUT benchmark dataset, demonstrating that our proposed CrackSCF network outperforms existing SOTA methods.
-The organization of the subsequent sections of this paper is as follows: In Section 2, existing crack detection models and lightweight networks are reviewed.
-Section 3 details the components of the proposed network.
-Section 4 delves into the experimental details and results.
-Finally, Section 5 provides a summary of the entire paper and highlights future research directions.
-2 Related Work
+2 Methodology
+2.1 Network Architecture
+As shown in Figure 2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a lightweight Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM).
+The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled.
+In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
 Figure 2: General Architecture Diagram of CrackSCF Network.
@@
 Through four stages, resulting in the segmentation output.
-2.1 Crack Segmentation Networks
-Recently, a growing number of researchers have utilized CNNs to develop crack segmentation networks, achieving advanced performance.
-Zhu et al.
-[13] proposed the BSNBM network, which integrates bidirectional self-rectifying modules with Bayesian inference to enhance crack detection robustness and accuracy under uncertainty.
-Cheng et al.
-[11] introduced the Selective Feature Fusion and Irregular-Aware Network (SFIAN) to effectively model irregular crack objects, resulting in clearer backgrounds on various crack images.
-Lyu et al.
-[14] proposed a CNN-based crack quantification framework that utilizes voxel-based reconstruction and Bayesian data fusion to achieve accurate crack width estimation.
-Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes.
-Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation.
-Qi et al.
-[15] developed an end-to-end model based on ViT and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders.
-Feng et al.
-[16] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
-Although CNN and CNN-Transformer methods have shown promise in crack segmentation, CNNs struggle to capture global context due to their local receptive fields, leading to discontinuous predictions and vulnerability to background noise.
-CNN-Transformer hybrids improve global perception but often adopt single-branch fusion, limiting their ability to integrate local and long-range cues.
-Moreover, both lack selective interaction and noise suppression, and their high computational cost hinders deployment on edge devices.
-2.2 Lightweight Networks
-Deploying deep learning models on edge devices like smartphones and drones is trending.
-However, these devices have limited computing resources, making it challenging to use models with high parameter counts.
-Thus, compact model design for lightweight modifications in size and speed is crucial.
-For example, Howard et al.
-[17] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices.
-Ma et al.
-[18] introduced ShuffleNet, providing practical design principles for efficient network architecture.
-Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures.
-Swaminathan et al.
-[19] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
-In recent years, there has been a growing trend among researchers to integrate model lightweighting techniques with crack segmentation models, aiming to facilitate real-time crack detection on edge computing devices.
-For example, Zhu et al.
-[20] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices.
-Huang et al.
-[21] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection.
-Although lightweight models reduce computational cost, their aggressive parameter reduction limits model expressiveness, impairing the capture of intricate patterns and leading to suboptimal segmentation.
-This trade-off between performance and efficiency undermines their reliability in real-world applications.
-3 Methodology
-3.1 Network Architecture
-As shown in Fig.
-2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a lightweight Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM).
-The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled.
-In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
-Specifically, for a given dataset, the images are first input into the MFE, which is divided into four stages.
+Specifically, for a given dataset, the crack images are first input into the MFE module, which is divided into four stages.
 In each stage, the extracted feature maps undergo a Feature Enhancement stage.
-Then, the LDE processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies.
-Subsequently, the SCFM receives these sequences along with the four layers of feature maps extracted by the MFE.
+Then, the LDE module processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies.
+Subsequently, the SCFM module receives these sequences along with the four layers of feature maps extracted by the MFE.
 After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
-Importantly, as shown in Fig.
-3, using the original convolution operations results in high FLOPs and Params, especially in the SCFM module, reaching 73.36G and 9.71M, respectively.
+Importantly, as shown in Figure 3, using the original convolution operations results in high FLOPs and Params, especially in the SCFM module, reaching 73.36G and 9.71M, respectively.
 Replacing all convolutions with LRDS blocks significantly reduced FLOPs and Params, demonstrating the effectiveness of LRDS in model lightweighting.
-Figure 3: Comparison of each module when using our LRDS or original convolution separately. (a) shows the comparison of FLOPs, (b) shows the comparison of parameters.
+Figure 3: Comparison of FLOPs and Params for each module using the original convolution and our proposed LRDS convolution block.
 2.2 Lightweight Multi-scale Feature Extractor
-As shown in Fig.
-2, our designed LRDS block adopts a bottleneck-like structure.
+As shown in Figure 4, our designed LRDS block adopts a bottleneck-like structure.
 It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer.
-The input feature map first passes through the Reduce layer, reducing the input channels to $f_{mid}$.
+The input feature map first passes through the Reduce layer, reducing the input channels to $C_{m}$.
 It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged.
-Finally, the Expand layer increases the number of channels to $f_{out}$, producing the final output feature map.
+Finally, the Expand layer increases the number of channels to $C_{d}$, producing the final output feature map.
 Low-rank approximation decomposes a high-dimensional matrix into low-dimensional matrices, reducing computation complexity.
 In CNNs, it is applied to convolutional kernel responses, assuming these lie in a low-rank subspace.
-Specifically, in the convolutional layer, we assume that its filter has a spatial size of $k$ and a number of input channels of $c$.
+Specifically, in the convolutional layer, the filter is assumed to have a spatial size of $k$ and a number of input channels of $c$.
 In computing the convolutional response, the filter is applied to a sub-tensor of the input tensor of shape $k\times k\times c$.
-Suppose this sub-tensor is $t\in\mathbb{R}^{k^{2}c+1}$, where we append a bias term at the end.
-The response at a position in the output, $y\in\mathbb{R}^{e}$, can be expressed as follows:
+Let this sub-tensor be denoted as $t\in\mathbb{R}^{k^{2}c+1}$, where a bias term is appended at the end.
+Consequently, the response at a position in the output, $y\in\mathbb{R}^{e}$, can be expressed as follows:
 $y=Wt$ (1)
-where $e$ represents the number of filters and $W$ is a $e\times(k^{2}c+1)$ matrix.
-Assuming the vector $y$ lies in a low-rank subspace, we can write it as $y=U(y-y_{1})+y_{1}$, where $U$ is a $e\times e$ matrix of rank $e_{0}$, with $e_{0}<e$.
-$y_{1}$ is the mean vector of the response.
-Therefore, we can write the response as
+where $e$ represents the number of filters and $W$ is an $e\times(k^{2}c+1)$ matrix.
+Assuming the vector $y$ lies in a low-rank subspace, it can be expressed as $y=U(y-y_{1})+y_{1}$, where $U$ is an $e\times e$ matrix of rank $e_{0}$, with $e_{0}<e$.
+Here, $y_{1}$ denotes the mean vector of the response.
+Therefore, the response can be formulated as
 $y=UWt+b$ (2)
 where $b=y_{1}-Uy_{1}$ is the new bias.
-The rank of $U$ is $e_{0}$, and it can be decomposed into two $e\times e_{0}$ matrices $J$ and $K$ such that $U=JK^{\top}$.
-We use $W_{1}=K^{\top}W$ to denote a $e_{0}\times(k^{2}c+1)$ matrix, which corresponds to a set of $e_{0}$ filters.
-Thus, we can rewrite the response as follows:
+Given that the rank of $U$ is $e_{0}$, it can be decomposed into two $e\times e_{0}$ matrices $J$ and $K$ such that $U=JK^{\top}$.
+Let $W_{1}=K^{\top}W$ denote an $e_{0}\times(k^{2}c+1)$ matrix, which corresponds to a set of $e_{0}$ filters.
+Thus, the response can be rewritten as follows:
 $y=JW_{1}t+b$ (3)
@@
 Given that $O(ee_{0})\ll O(e_{0}k^{2}c)$, the complexity of using Equation (3) is $\frac{e_{0}}{e}$ times that of using Equation (1).
-In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space.
-We constructed MFE by replacing all convolution operations with LRDS convolution blocks.
+Low-rank approximation theory significantly reduces the parameters and computational load of convolutional layers.
+To efficiently capture spatial information within the low-rank space for finer perception of detailed regions, the LRDS module processes features by embedding depthwise convolution and pointwise convolution directly into the low-rank space.
+Consequently, the MFE is constructed by replacing all standard convolution operations in the ResNet50 backbone with LRDS convolution blocks.
+Figure 4: Illustration of LRDS convolution block.
+After the number of channels of the input features is reduced, their local and spatial features are efficiently extracted by depth-wise convolution and point-wise convolution, and finally the number of channels is restored.
 2.3 Lightweight Long-range Dependency Extractor
@@
 Specifically, given an input $x\in\mathbb{R}^{C\times H\times W}$, let $q$ denote a query element with query feature $z_{q}$ and reference point coordinates $(p_{qx},p_{qy})$, the deformable attention is calculated as:
-$\centering\begin{split}&MSDeformAttn(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg[\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg]\end{split}\@add@centering$ (4)
+$\centering\begin{split}&\text{MSDeformAttn}(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg[\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg]\end{split}\@add@centering$ (4)
 Where $h$ represents the attention head, $c$ represents the input feature layer, and $t$ represents the sampling key point.
@@
 Here, normalized coordinates $\hat{p}\in[0,1]$ are used to represent the position of a reference point in the normalized feature map $\phi_{c}(\hat{p}_{q})$ rescales the normalized coordinates onto the $c$-th layer feature map.
-As shown in Fig.
-4(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module.
-In order to reduce the computational resources required by the LDE, we propose a linear transformation method (LR Linear) based on low-rank approximation.
+As shown in Figure 5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module.
+To reduce the computational resources required by the LDE, a linear transformation method (LR Linear) based on low-rank approximation is proposed.
 In traditional linear layers, the weight matrix connecting the input and output is typically a high-rank matrix, representing a complete connection between input and output channels.
-By applying low-rank approximation, we decompose the weight matrix into two smaller matrices, reducing the number of parameters.
+By applying low-rank approximation, the weight matrix is decomposed into two smaller matrices, effectively reducing the number of parameters.
 Specifically, the weight matrix $E\in\mathbb{R}^{r_{\text{out}}\times r_{\text{in}}}$, where $r_{\text{out}}$ is the output feature dimension and $r_{\text{in}}$ is the input feature dimension, is approximated by two smaller matrices $A\in\mathbb{R}^{r_{\text{out}}\times r}$ and $B\in\mathbb{R}^{r_{\text{in}}\times r}$, where $r$ is a small rank, much smaller than $r_{\text{out}}$ and $r_{\text{in}}$.
@@
 $E\approx E^{\prime}=AB^{T}$ (5)
-we perform this transformation through two linear layers.
-First, we apply a linear transformation $B^{T}$ to the input $x\in\mathbb{R}^{r_{\text{in}}}$, resulting in an intermediate representation $z\in\mathbb{R}^{r}$:
+This transformation is performed through two linear layers.
+First, a linear transformation $B^{\top}$ is applied to the input $x\in\mathbb{R}^{r_{\text{in}}}$, resulting in an intermediate representation $z\in\mathbb{R}^{r}$:
 $z=B^{T}x$ (6)
-Then, we apply another linear transformation $A$ to $z$, resulting in the output $y\in\mathbb{R}^{r_{\text{out}}}$:
+Subsequently, another linear transformation $A$ is applied to $z$, resulting in the output $y\in\mathbb{R}^{r_{\text{out}}}$:
 $y=Az$ (7)
@@
 This leads to a reduction in both the computational complexity and the number of parameters, thus improving training and inference efficiency.
-As shown in Fig.
-4(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
-The structure of LDE is shown in Fig.
-2.
+As shown in Figure 5(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
+The structure of LDE is shown in Figure 2.
 LDE facilitates long-range dependency modeling through the interaction of multi-scale features between the encoder and decoder.
@@
 The encoder then combines LR Linear with lightweight deformable attention mechanisms to capture long-range dependencies:
-$Z=MSDeformAttn(S,P,S)$ (8)
+$Z=\text{MSDeformAttn}(S,P,S)$ (8)
 where $P\in\mathbb{R}^{B\times L\times 2}$ represent the normalized reference coordinates.
 Subsequently, $Z$ is processed through a feedforward network (FFN) that combines LR Linear to refine the features:
-M=LR_Linear(GELU(LR_Linear(Z)))M=LR\_Linear(GELU(LR\_Linear(Z))) (9)
+$M=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Z)))$ (9)
 where the output memory matrix $M\in\mathbb{R}^{B\times L\times d}$ is used as the input to the decoder, encoding multi-scale local dependencies.
@@
 First, the reference points $H\in\mathbb{R}^{N\times 2}$ are mapped to the encoded memory $M$ to aggregate features:
-$Y_{0}=MSDeformAttn(Q,H,M)$ (10)
+$Y_{0}=\text{MSDeformAttn}(Q,H,M)$ (10)
 Then, it passes through the FFN to obtain the crack pixel sequence enriched with irregular long-range dependencies:
-$Y=LR\_Linear(GELU(LR\_Linear(Y_{0})))$ (11)
+$Y=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Y_{0})))$ (11)
 where Y is the output sequence of the LDE enriched with irregular long-range dependencies between pixels.
 2.4 Staircase Cascaded Fusion Module
-The design of our SCFM structure is shown in Fig.
-2.
-As a segmentation head, to enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we proposed the lightweight Pixel Attention Guided Fusion (PAF) module.
-The structure of PAF is shown in Fig.
-2.
+Figure 6: Illustration of the PAF. It effectively reduces required computing resources by embedding LRDS.
+The design of our SCFM structure is shown in Figure 2.
+As a segmentation head, to enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, this paper proposed the lightweight Pixel Attention Guided Fusion (PAF) module.
+The structure of PAF is shown in Figure 6.
 We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively.
 The output can be expressed as:
-$Output=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (12)
+$\sigma=\text{Sigmoid}(f_{c}(\vec{v_{c}})\cdot f_{d}(\vec{v_{d}}))$ (12)
+where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. Thus, the output of PAF can be represented as:
+$\text{Output}=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (13)
 where $\sigma$ indicates the probability that the two pixels belong to the same category.
@@
 To reduce computational resource usage, all convolution operations in this module utilize LRDS blocks.
-As seen in Fig.
-3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively.
+As seen in Figure 3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively.
 Thus, lightweighting these operations is crucial.
@@
 The processes of SCFM can be given by the following equation:
-$\begin{split}&F_{i}^{O}=C_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\right)\odot P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\end{split}$ (13)
+$\begin{split}&F_{i}^{\text{O}}=C_{\text{i}}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\right)\odot P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\end{split}$ (14)
 Where $F_{i}^{\text{O}}$ represents the output features of the $i$-th SCFM layer ($i$ ranges from 0 to 3), $F_{i}^{\text{MFE}}$ represents the $i$-th layer MFE features, $C_{\text{i}}$ is an operation chain: Concat followed by LRDS, ReLU, and GN, $P_{i}$ is a PAF operation, $\mu(\cdot)$ is the sigmoid function generating spatial attention weights, and $\odot$ represents the Hadamard product.
@@
 The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance.
-Previously, Binary Cross-Entropy (BCE) was often used, but it may not effectively distinguish the minority crack pixels in the image.
-Combining BCE with Dice loss addresses this issue.
-BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data [22] .
+Previously, Binary Cross-Entropy (BCE) [16] was often used, but it may not effectively distinguish the minority crack pixels in the image.
+Combining BCE with Dice loss [15] addresses this issue.
+BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data.
 Our loss $L$ can be calculated using Equation (15), which combines the above two loss functions for joint optimization.
@@
 where $\mathit{L_{B}}$ represents the BCE loss function and $\mathit{L_{D}}$ represents the Dice loss function.
-4 Experiments
-In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics.
-We then analyze results and model complexity, perform ablation experiments.
+In this section, the paper analyzes the differences between the TUT dataset and other public datasets, describes the experimental setup, and compares various methods and metrics.
+Subsequently, the results and model complexity are analyzed, and ablation experiments are performed.
+3 Experimental Results and Discussion
 3.1 Dataset Description
 Table 1: Comparison of Different Crack Datasets
-No. Datasets Year Num(train/val/test) Resolution Percentage of cracked pixels Scenarios
-1 CFD [23] 2016 118(76/11/31) 480 width and 320 height 1.60% bitumen
-2 GAPS509 [24] 2017 509(307/100/102) 540 width and 640 height 1.19% bitumen
-3 DeepCrack [12] 2019 537(300/100/247) 544 width and 384 height 3.54% bitumen, cement
-4 Crack500 [3] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen, cement
-5 CrackMap [25] 2023 120(72/12/36) 256 width and 256 height 4.10% bitumen
-6 TUT (Ours) 2024 1408(987/139/282) 640 width and 640 height 3.16% bitumen, cement, bricks, runways, metal, blades, pipelines, tiles
+No. Datasets Num Resolution Crack Pixels Scenarios
+1 CFD [28] 118 480 $\times$ 320 1.60% 1
+2 GAPS509 [9] 509 540 $\times$ 640 1.19% 1
+3 DeepCrack [17] 537 544 $\times$ 384 3.54% 2
+4 Crack500 [40] 3368 640 $\times$ 360 6.01% 2
+5 CrackMap [14] 120 256 $\times$ 256 4.10% 1
+6 AEL [40] 58 768 $\times$ 512 0.67% 1
+7 CrackTree206 [46] 206 800 $\times$ 600 0.32% 1
+8 LSCD [41] 143 400 $\times$ 400 3.18% 2
+9 CrackSC [12] 197 320 $\times$ 480 2.02% 2
+10 TUT (Ours) 1408 640 $\times$ 640 3.16% 8
+Figure 7: Illustration of the images of each scenarios in the TUT dataset.
 The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only.
 These datasets have clear backgrounds and minimal interference, which cannot comprehensively evaluate the performance of models.
-To tackle this problem, we created the TUT dataset to enable better segmentation across different scenarios.
+To tackle this, this paper created the TUT dataset to enable better segmentation across different scenarios.
 Detailed information and comparisons with other datasets are list in Table 1.
-The TUT dataset comprises 1270 crack images, along with a small number of 138 images sourced from the Internet.
+The TUT dataset comprises 1270 crack images captured with our mobile phones, along with a small number of 138 images sourced from the Internet.
 TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances.
 To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images.
-Unlike other datasets with clear backgrounds, the TUT dataset images have complex backgrounds and cracks of intricate shapes.
+Unlike other datasets with clear backgrounds, the TUT dataset images have complex, noisy backgrounds and cracks of intricate shapes.
 This evaluates the model’s capability to segment cracks in challenging conditions.
-Figure 5: Illustration of the various scenarios of the TUT dataset with the original images and Ground Truth.
-As list in Table 1, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios.
-Fig.
-5 shows crack images in different scenarios and their binary labeled images.
+As shown in Table 1, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios.
+Figure 7 shows crack images in different scenarios and their binary labeled images.
 The crack pixel ratio in the TUT dataset is 3.16%, which is moderate.
 This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.
-As seen in Fig.
-5, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties.
+As seen in Figure 7, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties.
 Particularly in plastic runways, both coarse and fine cracks are present, testing the model’s generalization to different types of cracks.
@@
 We compared the CrackSCF network with other SOTA methods on six datasets: CFD [28], GAPS509 [9], DeepCrack [17], Crack500 [40], CrackMap [14], and ours TUT dataset.
-Table 2: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
-Methods CFD GAPS509 Crack500
-ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU
-RIND 0.6858 0.7021 0.6456 0.7534 0.6954 0.7592 0.5505 0.6023 0.4959 0.6494 0.5623 0.6909 0.6469 0.6483 0.6998 0.7245 0.7119 0.7381
-UCTNet 0.6587 0.7006 0.6701 0.7205 0.6944 0.7479 0.5835 0.6188 0.5709 0.6752 0.6187 0.7107 0.7147 0.7265 0.6887 0.7843 0.7302 0.7674
-SFIAN 0.7040 0.7159 0.6657 0.7512 0.7058 0.7684 0.6087 0.6320 0.6061 0.6451 0.6250 0.7198 0.6977 0.7348 0.6983 0.7742 0.7343 0.7604
-CTCS 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
-DTrCNet 0.7079 0.7123 0.6774 0.7486 0.7112 0.7710 0.6211 0.6276 0.6277 0.6592 0.6431 0.7272 0.7012 0.7241 0.6527 0.8280 0.7357 0.7627
-Crackmer 0.6751 0.6754 0.6370 0.7237 0.6776 0.7503 0.4593 0.4642 0.6144 0.4663 0.5302 0.6581 0.6933 0.7097 0.6985 0.7572 0.7267 0.7591
-SimCrack 0.7152 0.7204 0.6883 0.7557 0.7204 0.7756 0.6408 0.6471 0.6171 0.6997 0.6558 0.7368 0.7127 0.7308 0.7093 0.7984 0.7516 0.7715
-Ours 0.7335 0.7376 0.6967 0.7848 0.7383 0.7864 0.6492 0.6552 0.6459 0.6812 0.6631 0.7411 0.7226 0.7353 0.7213 0.7922 0.7551 0.7764
-CrackMap DeepCrack TUT (Ours)
-RIND 0.6745 0.6943 0.6023 0.7586 0.6699 0.7425 0.8087 0.8267 0.7896 0.8920 0.8377 0.8391 0.7531 0.7891 0.7872 0.7665 0.7767 0.8051
-UCTNet 0.6104 0.6195 0.5653 0.6748 0.6152 0.7079 0.8357 0.8579 0.8217 0.8857 0.8525 0.8564 0.7569 0.7670 0.8201 0.7667 0.7925 0.8110
-SFIAN 0.7200 0.7465 0.6715 0.7668 0.7160 0.7748 0.8616 0.8928 0.8549 0.8692 0.8620 0.8776 0.7290 0.7513 0.7715 0.7367 0.7537 0.7896
-CTCS 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
-DTrCNet 0.7328 0.7413 0.6912 0.7681 0.7276 0.7812 0.8473 0.8512 0.8905 0.8251 0.8566 0.8661 0.7987 0.8073 0.7972 0.8441 0.8202 0.8331
-Crackmer 0.7395 0.7437 0.7229 0.7467 0.7346 0.7860 0.8712 0.8785 0.8946 0.8783 0.8864 0.8844 0.7429 0.7640 0.7501 0.7656 0.7578 0.7966
-SimCrack 0.7559 0.7625 0.7380 0.7672 0.7523 0.7963 0.8570 0.8722 0.8984 0.8549 0.8761 0.8744 0.7984 0.8090 0.8051 0.8371 0.8208 0.8334
-Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
-2) Comparison Methods: To evaluate the model performance, we compared the CrackSCF with seven classical and latest SOTA methods on different datasets.
-These methods include RIND [26], UCTNet [8], SFIAN [11], CTCS [9], DTrCNet [7], Crackmer [10] and Simcrack [27].
-3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics [11]: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
+2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation methods and the latest SOTA methods on different datasets.
+These methods include HED [37], UNet++ [44], Deeplabv3+ [2], AttuNet [22], RCF [19], RIND [24], UCTNet [33], SFIAN [4], CT-crackseg [30], DTrCNet [35], Crackmer [34] and Simcrack [13].
+Figure 8: Visualization results of CrackSCF and other methods on public datasets.
+Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
+3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), Mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
+Mean Intersection over Union (mIoU) measures the average ratio of the intersection and union of the predicted results and the true results.
+The calculation formula for mIoU is:
+$mIoU=\frac{1}{i+1}\sum_{a=0}^{i}\frac{p_{aa}}{\sum_{b=0}^{i}p_{ab}+\sum_{b=0}^{i}p_{ba}-p_{aa}}$ (16)
+where $i$ is the number of classes, $a$ denotes the true value, $b$ denotes the predicted value, and $p_{ab}$ represents the number of pixels predicted as $a$ for $b$.
+Given that our crack segmentation task involves binary classification, we set $i=1$.
+Optimal Dataset Scale (ODS) indicates the optimal segmentation performance metric obtained when a fixed threshold $t$ is selected over the entire dataset. It is defined as:
+$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (17)
+Optimal Image Scale (OIS) indicates the optimal segmentation performance metric obtained when an optimal threshold $t$ is selected for each image individually. It is defined as:
+$OIS=\frac{1}{N}\sum_{a=1}^{N}\max_{t}\frac{2\cdot P_{t,a}\cdot R_{t,a}}{P_{t,a}+R_{t,a}}$ (18)
+where $N$ represents the total number of images.
 3.3 Comparison with the SOTA methods
-1) Results on the our TUT dataset: Fig.
-6 shows the visual results of all methods on the TUT dataset.
-We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes.
-As list in Table 2, our method achieved excellent results on the TUT dataset, leading in all metrics.
+Our proposed method, CrackSCF, was compared with 12 other methods across five public datasets and our own TUT dataset.
+Table 2: Comparison results on the CFD dataset. Best results are bolded and second best results are underlined.
+Methods ODS OIS P R F1 mIoU
+HED 0.6276 0.6494 0.5467 0.7374 0.6279 0.7240
+UNet++ 0.6586 0.6836 0.5995 0.7608 0.6706 0.7432
+DeeplabV3+ 0.6910 0.7084 0.6264 0.7730 0.6920 0.7709
+AttuNet 0.6710 0.6870 0.6240 0.7220 0.6694 0.7576
+RCF 0.6434 0.6624 0.5792 0.7348 0.6478 0.7343
+RIND 0.6858 0.7021 0.6456 0.7534 0.6954 0.7592
+UCTNet 0.6587 0.7006 0.6701 0.7205 0.6944 0.7479
+SFIAN 0.7040 0.7159 0.6657 0.7512 0.7058 0.7684
+CT-crackseg 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738
+DTrCNet 0.7079 0.7123 0.6774 0.7486 0.7112 0.7710
+Crackmer 0.6751 0.6754 0.6370 0.7237 0.6776 0.7503
+SimCrack 0.7152 0.7204 0.6883 0.7557 0.7204 0.7756
+Ours 0.7335 0.7376 0.6967 0.7848 0.7383 0.7864
+1) Results on the CFD dataset: The crack areas in the CFD dataset are very slender, testing the model’s sensitivity to fine-grained crack pixels.
+As shown in Table 2, our proposed CrackSCF attained the best results across all metrics for the CFD dataset.
+Specifically, on ODS, OIS, P, R, F1, and mIoU, it outperformed the second-best methods by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively.
+Notably, SimCrack [13] achieved the second-best performance on all metrics except for P and R, only behind our method.
+However, as seen in Figure 8, its actual segmentation performance is not satisfactory, while our method achieves high-quality segmentation results on various fine-grained crack pixels.
+Other methods, such as UCTNet [33] and DTrCNet [35], tend to misdetect background noise or miss key detail areas of crack pixels.
+Table 3: Comparison results on the GAPS509 dataset. Best results are bolded and second best results are underlined.
+Methods ODS OIS P R F1 mIoU
+HED 0.4014 0.4509 0.4300 0.4624 0.4456 0.6282
+UNet++ 0.6255 0.6484 0.6160 0.6684 0.6411 0.7300
+DeeplabV3+ 0.6160 0.6425 0.6050 0.6669 0.6345 0.7378
+AttuNet 0.5955 0.6174 0.5924 0.6519 0.6208 0.7247
+RCF 0.4532 0.4817 0.4420 0.5428 0.4872 0.6472
+RIND 0.5505 0.6023 0.4959 0.6494 0.5623 0.6909
+UCTNet 0.5835 0.6188 0.5709 0.6752 0.6187 0.7107
+SFIAN 0.6087 0.6320 0.6061 0.6451 0.6250 0.7198
+CT-crackseg 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165
+DTrCNet 0.6211 0.6276 0.6277 0.6592 0.6431 0.7272
+Crackmer 0.4593 0.4642 0.6144 0.4663 0.5302 0.6581
+SimCrack 0.6408 0.6471 0.6171 0.6997 0.6558 0.7368
+Ours 0.6492 0.6552 0.6459 0.6812 0.6631 0.7411
+2) Results on the GAPS509 dataset: Unlike the CFD dataset, the GAPS509 dataset contains not only finer cracks but also some irrelevant background interference, testing the model’s ability to extract crack areas under high interference conditions.
+As shown in Table 3, our method achieved the highest performance on all metrics except R.
+Although SimCrack [13] achieved 0.6997 on the R metric, leading our method by about 2.72%, its P metric was 0.6171, far lower than our method.
+The highest F1 score obtained by our method indicates a good balance between P and R.
+Specifically, our method outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.31%, 1.25%, 2.90%, 1.11%, and 0.45%, respectively.
+Additionally, as seen in Figure 8, our method almost perfectly avoids irrelevant background noise, while most other methods falsely detect irrelevant backgrounds and perform worse in segmenting crack pixel areas compared to our method.
+Table 4: Comparison results on the DeepCrack dataset. Best results are bolded and second best results are underlined.
+Methods ODS OIS P R F1 mIoU
+HED 0.7552 0.8237 0.7231 0.7678 0.7448 0.8007
+UNet++ 0.8419 0.8893 0.8474 0.8533 0.8503 0.8621
+DeeplabV3+ 0.7727 0.8354 0.7363 0.8323 0.7813 0.8565
+AttuNet 0.8153 0.8379 0.8066 0.8069 0.8068 0.8578
+RCF 0.7666 0.8376 0.7423 0.7611 0.7516 0.8110
+RIND 0.8087 0.8267 0.7896 0.8920 0.8377 0.8391
+UCTNet 0.8357 0.8579 0.8217 0.8857 0.8525 0.8564
+SFIAN 0.8616 0.8928 0.8549 0.8692 0.8620 0.8776
+CT-crackseg 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925
+DTrCNet 0.8473 0.8512 0.8905 0.8251 0.8566 0.8661
+Crackmer 0.8712 0.8785 0.8946 0.8783 0.8864 0.8844
+SimCrack 0.8570 0.8722 0.8984 0.8549 0.8761 0.8744
+Ours 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001
+3) Results on the DeepCrack dataset: The cracks in the DeepCrack dataset are often relatively wide, testing the model’s ability to extract a large range of crack pixels.
+As shown in Table 4, our method achieved the highest results on all metrics for the DeepCrack dataset.
+Specifically, our method outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85%, respectively.
+Notably, CT-crackseg also achieved good results on this dataset, but as seen in Figure 8, its segmentation performance on detailed cracks is not as good as our method, indicating that our method can effectively handle crack details.
+4) Results on the Crack500 dataset: Similar to the DeepCrack, the Crack500 dataset also has relatively wide crack areas, but its structure is complex, posing higher requirements on model performance.
+As shown in Table 5, our method achieved the best results on five metrics for the Crack500 dataset, except for R.
+It outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%, respectively.
+Although DTrCNet [35] leads on the R metric, our method surpasses it on the P metric by 10.51%.
+The highest F1 score achieved by our method indicates a good balance between P and R.
+Additionally, as seen in Figure 8, our method does not produce false detections in irrelevant background areas, whereas other methods either produce false detections or lack detailed segmentation in crack areas.
+Table 5: Comparison results on the Crack500 dataset. Best results are bolded and second best results are underlined.
+Methods ODS OIS P R F1 mIoU
+HED 0.6191 0.6537 0.6150 0.7281 0.6668 0.7140
+UNet++ 0.6965 0.7274 0.6989 0.7805 0.7374 0.7605
+DeeplabV3+ 0.6943 0.7211 0.6970 0.7569 0.7257 0.7746
+AttuNet 0.6586 0.6890 0.6490 0.7608 0.7005 0.7543
+RCF 0.6054 0.6361 0.6037 0.7136 0.6540 0.7068
+RIND 0.6469 0.6483 0.6998 0.7245 0.7119 0.7381
+UCTNet 0.7147 0.7265 0.6887 0.7843 0.7302 0.7674
+SFIAN 0.6977 0.7348 0.6983 0.7742 0.7343 0.7604
+CT-crackseg 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
+DTrCNet 0.7012 0.7241 0.6527 0.8280 0.7357 0.7627
+Crackmer 0.6933 0.7097 0.6985 0.7572 0.7267 0.7591
+SimCrack 0.7127 0.7308 0.7093 0.7984 0.7516 0.7715
+Ours 0.7226 0.7353 0.7213 0.7922 0.7551 0.7764
+5) Results on the CrackMap dataset: Similar to the CFD dataset, the cracks in the CrackMap dataset are also relatively slender.
+As shown in Table 6, our method leads all other methods on all metrics, outperforming the second-best methods on ODS, OIS, P, R, F1, and mIoU by 0.48%, 0.34%, 0.14%, 0.35%, 0.31%, and 0.30%, respectively.
+Notably, SimCrack [13] achieved the second-best results on this dataset, but as seen in Figure 8, SimCrack’s segmentation performance on the crack tips is not satisfactory, whereas our method effectively identifies these pixels.
+Table 6: Comparison results on the CrackMap dataset. Best results are bolded and second best results are underlined.
+Methods ODS OIS P R F1 mIoU
+HED 0.6022 0.6240 0.5281 0.7063 0.6043 0.7026
+UNet++ 0.7030 0.7222 0.6773 0.7209 0.6984 0.7631
+DeeplabV3+ 0.6617 0.6793 0.5853 0.7532 0.6587 0.7460
+AttuNet 0.6618 0.6773 0.6233 0.6970 0.6581 0.7454
+RCF 0.5552 0.5683 0.4590 0.7164 0.5595 0.6758
+RIND 0.6745 0.6943 0.6023 0.7586 0.6699 0.7425
+UCTNet 0.6104 0.6195 0.5653 0.6748 0.6152 0.7079
+SFIAN 0.7200 0.7465 0.6715 0.7668 0.7160 0.7748
+CT-crackseg 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785
+DTrCNet 0.7328 0.7413 0.6912 0.7681 0.7276 0.7812
+Crackmer 0.7395 0.7437 0.7229 0.7467 0.7346 0.7860
+SimCrack 0.7559 0.7625 0.7380 0.7672 0.7523 0.7963
+Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987
+Table 7: Comparison results on the TUT dataset. Best results are bolded and second best results are underlined.
+Methods ODS OIS P R F1 mIoU
+HED 0.6960 0.7254 0.7080 0.7431 0.7251 0.7683
+UNet++ 0.7702 0.7948 0.7725 0.7974 0.7848 0.8160
+DeeplabV3+ 0.7504 0.7777 0.7155 0.8231 0.7655 0.8177
+AttuNet 0.7539 0.7795 0.7558 0.7945 0.7747 0.8204
+RCF 0.6821 0.7138 0.7106 0.6877 0.6990 0.7598
+RIND 0.7531 0.7891 0.7872 0.7665 0.7767 0.8051
+UCTNet 0.7569 0.7670 0.8201 0.7667 0.7925 0.8110
+SFIAN 0.7290 0.7513 0.7715 0.7367 0.7537 0.7896
+CT-crackseg 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
+DTrCNet 0.7987 0.8073 0.7972 0.8441 0.8202 0.8331
+Crackmer 0.7429 0.7640 0.7501 0.7656 0.7578 0.7966
+SimCrack 0.7984 0.8090 0.8051 0.8371 0.8208 0.8334
+Ours 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
+Figure 9: Visualization results of CrackSCF and other methods on TUT dataset.
+Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
+6) Results on the TUT dataset: Figure 9 presents the visual results of all methods on the TUT dataset.
+Easily missed crack pixel areas are highlighted with red boxes, and misdetected non-crack pixel areas are indicated with yellow boxes.
+As shown in Table 7, the proposed method achieved excellent results on the TUT dataset, leading in all metrics.
 Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively.
-SimCrack [27] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig.
-6, its segmentation performance on some key details has room for improvement compared to our method.
+While SimCrack [13] achieved the second-best results on ODS, OIS, F1, and mIoU, as observed in Figure 9, its segmentation performance on certain key details exhibits room for improvement compared to the proposed method.
 Furthermore, most methods fail to suppress irrelevant background noise.
@@
 Our method demonstrates excellent robustness and versatility, resulting in high-quality crack segmentation maps and adapting well to different scenarios.
-2) Results on other public datasets: The CFD, CrackMap and GAPS509 datasets are characterized by slender and fine-grained crack structures, testing the model’s sensitivity to subtle and narrow crack pixels.
-As listed in Table 2, our proposed CrackSCF achieved the best performance on the CFD and CrackMap datasets.
-On both datasets, CrackSCF outperforms Simcrack by an average of 1.97%, 9.89%, 2.85%, 1.83%, and 1.00% on ODS, OIS, P, F1, and mIoU respectively.
-Although SimCrack [27] ranked second on most metrics, its actual segmentation performance on fine crack pixels and crack tips was unsatisfactory, as shown in Fig.
-6.
-Other methods such as UCTNet [8] and DTrCNet [7] often failed to distinguish crack regions from background noise or missed crucial fine-grained crack details.
-On the more challenging GAPS509 dataset, our method also achieved the highest performance on all metrics except recall.
-SimCrack obtained a higher recall value of 0.6997, about 2.72% higher than our method, but its precision was significantly lower at 0.6171, leading to a lower F1 score overall.
-Our method maintained a superior balance between precision and recall, with improvements of 1.31%, 1.25%, 2.90%, 1.11%, and 0.45% on ODS, OIS, P, F1, and mIoU, respectively.
-Visual results in Fig.
-6 further confirm that our method effectively suppresses false detections in irrelevant background areas while preserving accurate segmentation of slender cracks.
-Figure 6: Visualization results of CrackSCF and other methods on TUT and other public dataset.
-Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
-Both DeepCrack and Crack500 datasets are characterized by relatively wide crack regions, testing the model’s capability to capture extensive crack areas.
-As listed in Table 2, our method achieved the best results across all metrics on the DeepCrack dataset, outperforming the second-best methods by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85% in terms of ODS, OIS, P, R, F1, and mIoU.
-Although CTCS [9] also performed well, it fell short in preserving crack detail segmentation, as illustrated in Fig.
-6.
-This demonstrates the advantage of our method in accurately segmenting both wide crack regions and their finer structures.
-On the Crack500 dataset, our method achieved leading performance in five of the six evaluation metrics: ODS, OIS, P, F1, and mIoU, surpassing the second-best methods by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%.
-While DTrCNet [7] achieved a higher recall, our method outperformed it in precision by a notable 10.51%, leading to the highest F1 score and indicating a superior balance between detecting true crack pixels and suppressing false positives.
-Furthermore, as shown in Fig.
-6, our method produces fewer erroneous detections in irrelevant background areas and achieves more accurate and refined segmentation of complex crack structures compared to competing approaches.
-Table 3: Comparison of the complexity of CrackSCF with other methods
+Table 8: Comparison of the complexity of CrackSCF with other methods. Best results are bolded and second best results are underlined.
 Methods FLOPs ↓ Params ↓ FPS ↑
+HED 80.35G 14.72M 33
+RCF 115.54G 15.52M 30
 RIND 695.77G 59.39M 11
 UCTNet 171.77G 66.22M 12
+UNet++ 139.61G 9.16M 22
+DeeplabV3+ 88.53G 59.34M 29
+AttuNet 541.34G 57.16M 24
 SFIAN 52.36G 13.63M 35
-CTCS 39.47G 22.88M 28
+CT-crackseg 39.47G 22.88M 28
 DTrCNet 123.20G 63.45M 47
+SimCrack 70.66G 29.58M 29
 Crackmer 14.94G 5.90M 31
-SimCrack 70.66G 29.58M 29
 Ours 9.26G 4.79M 36
-Table 4: Comparison of CrackSCF using LRDS blocks with networks using other convolutions
-Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
-Original Conv 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
-DS Conv 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
-Low-Rank Conv 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
-LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
-Table 5: Comparison of SCFM performance with other segmentation heads
-SegHead ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
-PID 0.7120 0.7198 0.7921 0.6742 0.7284 0.7777 36.69G 5.44M 21
-SegFormer 0.7446 0.7634 0.7533 0.7786 0.7658 0.7983 29.72G 6.20M 23
-UNet 0.7836 0.7951 0.7829 0.8248 0.8033 0.8222 40.59G 12.34M 20
-Hamburger 0.7447 0.7609 0.7075 0.8328 0.7670 0.7969 168.23G 7.45M 14
-SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
-Table 6: Comparison of the performance of using our lightweight deformable attention in LDE with other variants
-Attention ODS OIS P R F1 mIoU FLOPs Params FPS↑
-Common MHA 0.7919 0.8009 0.7999 0.8240 0.8118 0.8279 15.98G 5.80M 16
-MHA + LR_Linear 0.7924 0.8003 0.7988 0.8284 0.8134 0.8286 8.87G 4.77M 14
-Common MSDA 0.8058 0.8151 0.8052 0.8467 0.8257 0.8373 16.37G 5.82M 28
-MSDA + LR_Linear 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
 3.4 Complexity Analysis
-To evaluate our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS).
+To evaluate the efficacy of the LRDS block, its computational and parameter efficiency was analyzed using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS).
 FLOPs quantify computational load, reflecting demand during inference.
@@
 Table 8 presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS.
-Our approach achieves notably low values in FLOPs at 9.26G and Params at 4.79M, surpassing the second-best by 38.01% and 18.81%, respectively.
+Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively.
 This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices.
@@
 3.5 Ablation Studies
-We conducted ablation studies on the TUT dataset to assess the effectiveness of our proposed module.
-1) Ablation study of different convolution types: As list in Table 4, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution.
-Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS.
+To evaluate the strengths of the proposed module over other modules, ablation experiments were performed on the TUT dataset.
+Table 9: Comparison of performance metrics using different convolution blocks. Best results are bolded and second best results are underlined.
+Method ODS OIS P R F1 mIoU
+Original Conv 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437
+DS Conv 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417
+Low-Rank Conv 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441
+LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
+Table 10: Comparison of computational efficiency. Best results are bolded and second best results are underlined.
+Method FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
+Original Conv 90.06G 17.86M 23
+DS Conv 23.91G 8.38M 35
+Low-Rank Conv 21.38G 4.63M 30
+LRDS Conv 9.26G 4.79M 36
+1) Ablation study of different convolution types: As shown in Table 9, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution.
+As shown in Table 10, Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS.
 Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance.
 These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
-2) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [28], SegFormer [6], UNet [29], and Hamburger head [30], using the TUT dataset.
-As list in the Table 5, SCFM achieves the best results in segmentation performance metrics.
+Table 11: Comparison of segmentation performance using SCFM, Single Concat branch, and Single PAF branch.
+Method ODS OIS P R F1 mIoU
+Single Concat branch 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359
+Single PAF branch 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427
+SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473
+Table 12: Comparison of computational efficiency. Best results are bolded and second best results are underlined.
+Method FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
+Single Concat branch 7.21G 4.71M 42
+Single PAF branch 8.24G 4.72M 39
+SCFM 9.26G 4.79M 36
+2) Ablation study in SCFM using different branch: To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, ablation experiments were performed on the TUT dataset, As shown in Table 11 and 12.
+Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics.
+Using only the PAF branch results in a 1.06% drop in the F1 score.
+While SCFM slightly increases parameter count, computational load, and inference speed, it significantly improves segmentation performance.
+These results indicate that our module effectively balances performance and efficiency.
+Table 13: Comparison of segmentation performance of SCFM with other segmentation heads. Best results are bolded and second best results are underlined.
+SegHead ODS OIS P R F1 mIoU
+PID 0.7120 0.7198 0.7921 0.6742 0.7284 0.7777
+SegFormer 0.7446 0.7634 0.7533 0.7786 0.7658 0.7983
+UNet 0.7836 0.7951 0.7829 0.8248 0.8033 0.8222
+Hamburger 0.7447 0.7609 0.7075 0.8328 0.7670 0.7969
+SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473
+Table 14: Comparison of computational efficiency of SCFM with other segmentation heads. Best results are bolded and second best results are underlined.
+SegHead FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
+PID 36.69G 5.44M 21
+SegFormer 29.72G 6.20M 23
+UNet 40.59G 12.34M 20
+Hamburger 168.23G 7.45M 14
+SCFM 9.26G 4.79M 36
+3) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, ablation experiments were conducted on different segmentation heads, including PID [38], SegFormer [36], UNet [27], and Hamburger head [11], using the TUT dataset.
+As shown in the Table 13, SCFM achieves the best results in segmentation performance metrics.
 Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively.
 This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels.
-Furthermore, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS.
+Furthermore, as shown in 14, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS.
 This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
-3) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA.
-As shown in the Table 6, our method achieves the best results in segmentation performance metrics.
-Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively.
-Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of our use of MSDA.
+Table 15: Comparison of segmentation performance using our lightweight deformable attention in LDE with other variants. Best results are bolded and second best results are underlined.
+Attention ODS OIS P R F1 mIoU
+Common MHA 0.7919 0.8009 0.7999 0.8240 0.8118 0.8279
+MHA + LR 0.7924 0.8003 0.7988 0.8284 0.8134 0.8286
+Common MSDA 0.8058 0.8151 0.8052 0.8467 0.8257 0.8373
+MSDA 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473
+Table 16: Comparison of computational efficiency with different attention variants. Best results are bolded and second best results are underlined.
+Attention FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
+Common MHA 15.98G 5.80M 16
+MHA + LR 8.87G 4.77M 14
+Common MSDA 16.37G 5.82M 28
+MSDA + LR 9.26G 4.79M 36
+4) Ablation study of different attention variants in LDE: To validate that the LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, lightweight MSDA was compared with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA.
+As shown in Table 15 and 16, the proposed method achieves the best results in segmentation performance metrics.
+Specifically, compared to the common MSDA, the proposed method improves by 1.51% and 1.19% in F1 and mIoU, respectively.
+Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of utilizing MSDA.
 Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease.
 For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively.
-This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
+This demonstrates that the LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
+Table 17: Comparison of results of CrackSCF networks trained with different loss ratios. Best results are bolded and second best results are underlined
+Dice : BCE ODS OIS P R F1 mIoU
+1 : 5 0.8129 0.8203 0.8217 0.8458 0.8336 0.8424
+1 : 4 0.8049 0.8145 0.8116 0.8365 0.8239 0.8368
+1 : 3 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
+1 : 2 0.8146 0.8211 0.8227 0.8435 0.8330 0.8438
+1 : 1 0.8136 0.8204 0.8231 0.8436 0.8332 0.8431
+Only BCE 0.8003 0.8056 0.8349 0.8004 0.8173 0.8332
+Only Dice 0.8107 0.8165 0.8302 0.8265 0.8283 0.8409
+5) Analysis of different hyperparameter settings in the loss function: Table 17 shows the results of ablation experiments on different weight proportions of the two sub-losses in the loss function.
+As shown, a Dice loss to BCE loss ratio of 1:3 yields the highest values in five metrics: ODS, OIS, R, F1, and mIoU.
+Although the P metric is 0.80% lower than using only BCE loss, the F1 metric improves by 2.58%, indicating a better balance between P and R.
+Therefore, this ratio is adopted during training.
+In summary, a Dice loss to BCE loss ratio of 1:3 not only enhances pixel-level performance but also improves the global segmentation of crack areas, achieving superior segmentation performance.
+3.6 Limitations and Discussions
+Through these experiments, it is evident that SCFM effectively perceives and fuses local details and long-range pixel dependencies.
+The proposed network excels in various metrics on public datasets with simple scenes and performs exceptionally on the complex TUT dataset.
+The lightweight LRDS block minimizes parameters and computational load, facilitating deployment on resource-constrained devices like drones and smartphones.
+Additionally, LDE is efficient in capturing irregular long-distance dependencies between cracked pixels.
+However, the proposed network presents several limitations, as follows:
+Figure 10: Examples of CrackSCF misdetections and misses. The critical regions are marked with blue boxes.
+  1. 1.
+    Although our method generally achieves good crack segmentation with effective background noise suppression, as shown in the Figure 10, our model occasionally misdetects noise regions that resemble cracks in some images.
+It also fails to detect watermark occlusions with colors similar to the background.
+Therefore, it is essential to enhance modules such as MFE and LDE to improve the model’s capability to perceive and extract features from critical noise and occlusion areas, thereby further increasing its robustness and stability in complex scenarios.
+  2. 2.
+    Despite our method achieves the minimum parameter count and computational load, there is still room for improvement in inference speed.
+For example, our network achieves a processing speed of 36 frames per second, but the DTrCNet [35], which focuses more on inference speed, attains a faster speed of 47 frames per second.
+Therefore, it is necessary to further simplify convolution operations or the network’s complexity to achieve faster inference speeds while maintaining a good balance between segmentation performance and model lightweighting.
 4 Conclusion
 This paper proposes a staircase cascaded fusion crack segmentation network (CrackSCF) for pixel-level crack detection.
-In CrackSCF, we propose the lightweight convolutional block LRDS to construct the MFE, which dramatically reduces the amount of computation and parametric quantities required by the model and is able to efficiently capture the local details of the crack image, and we propose LDE to more efficiently perceive the pixel irregular long-range dependencies in the crack feature maps, in order to more comprehensively fuse the local patterns with the pixel’s long-range dependencies, we also propose SCFM to fully perceive them to generate higher quality segmentation maps.
-Furthermore, to comprehensively evaluate our model’s robustness and stability across various complex scenarios, we collected the TUT dataset, which includes eight types of crack scenes.
-Experimental results show that our method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation.
+Within CrackSCF, the lightweight convolutional block LRDS is introduced to construct the MFE.
+This block dramatically reduces the computational load and parameter quantity required by the model while efficiently capturing the local details of crack images.
+Additionally, the LDE is proposed to more efficiently perceive irregular long-range dependencies among pixels in the crack feature maps.
+To comprehensively fuse local patterns with these long-range dependencies, the SCFM is employed to generate higher-quality segmentation maps.
+Furthermore, to comprehensively evaluate the model’s robustness and stability across various complex scenarios, the TUT dataset, which includes eight types of crack scenes, was collected.
+Experimental results show that the proposed method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation.
 The computational load and parameter count are reduced to 9.26G and 4.79M, respectively.
-Despite our method achieves the low parameter count and computational load, there is still room for improvement in inference speed.
-In future work, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications.
-Additionally, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.
+Although the proposed method achieves a low parameter count and computational load, there is still room for improvement in inference speed.
+In future work, network lightweighting will be further pursued to reduce computational resource requirements and achieve higher inference speeds, with the aim of deploying the model on mobile devices for enhanced real-world applications.
+Additionally, future efforts will focus on designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.